### v4

# Sections of Lagrangian fibrations on holomorphic symplectic manifolds

Abstract. Let $M$ be a holomorphically symplectic manifold, equipped with a Lagrangian fibration $\pi:\;M\to X$. A degenerate twistor deformation (a more general version of it is called “a Tate-Shafarevich twist”) is a family of holomorphically symplectic structures on $M$ parametrized by $H^{1,1}(X)$. All members of this family are equipped with a holomorphic Lagrangian projection to $X$, and their fibers are isomorphic to the fibers of $\pi$. Assume that $M$ is a compact hyperkähler manifold of maximal holonomy, and the general fiber of the Lagrangian projection $\pi$ is primitive (that is, not divisible) in integer homology. We also assume that $\pi$ has reduced fibers in codimension 1. Then $M$ has a degenerate twistor deformation $M^{\prime}$ such that the Lagrangian projection $\pi:\;M^{\prime}\to X$ admits a meromorphic section.
1 Introduction
1.1 Holomorphic section of Lagrangian fibrations
A hyperkähler manifold, for our present purposes, is compact, Kähler, holomorphically symplectic manifold. We say that $M$ is a hyperkähler manifold of maximal holonomy, or IHS (irreducible holomorphically symplectic) if $H^{2,0}(M)={\mathbb{C}}$ and $\pi_{1}(M)=0$. The compactness and the maximal holonomy condition is tacitly assumed throughout this paper.
A Lagrangian fibration on a hyperkähler manifold is a holomorphic map $\pi:\;M{\>\longrightarrow\>}X$, with all fibers Lagrangian with respect to the holomorphic symplectic form on $M$. The manifold $X$ is biholomorphic to ${\mathbb{C}}P^{n}$ if it is smooth ([Hw]). With each Lagrangian fibration $\pi:\;M{\>\longrightarrow\>}X$ we associate a deformation $I_{t}$ of a complex structure on $M$, parametrized by $H^{2}(X,{\mathbb{C}})$ ([V1]). This deformation, called the degenerate twistor deformation, is properly introduced in Subsection 2.2. It shares many properties with the twistor deformation, known from hyperkähler geometry; in particular, its general fiber is non-algebraic. Most importantly, each member of this family is equipped with a holomorphic Lagrangian projection to $X$, and the fibers of this projection are isomorphic to the fibers of $\pi$. The main result of the present paper is the following theorem.
Theorem 1.1:  Let $\pi:\;M{\>\longrightarrow\>}X$ be a Lagrangian fibration on a compact hyperkähler manifold of maximal holonomy, with $X={\mathbb{C}}P^{n}$. Assume that the general fiber of the Lagrangian projection $\pi$ is primitive (that is, not divisible) in integer homology. Assume, moreover, that the fibers of $\pi$ outside of a codimension 2 subset in $X$ are reduced. Denote by $(M,I_{t})$ the degenerate twistor deformation, with $t\in H^{2}(X,{\mathbb{C}})$. Then for some $t_{0}\in H^{2}(X,{\mathbb{C}})$, the Lagrangian projection $\pi:\;(M,I_{t_{0}}){\>\longrightarrow\>}X$ admits a meromorphic section.
Proof: 5.3.
1.2 C-symplectic structures and degenerate twistor deformations
A superficially similar result was proven in [BDV]. Recall that a closed complex-valued 2-form $\Omega$ on a smooth manifold $M$, $\dim_{\mathbb{R}}M=4n$, is called C-symplectic if $\Omega^{n}\wedge\overline{\Omega}^{n}$ is non-degenerate, and $\Omega^{n+1}=0$. By [BDV, Proposition 2.13], for any C-symplectic form $\Omega$ on $M$ there exists a unique complex structure operator $I$ such that $\Omega$ is a holomorphic symplectic form on $(M,I)$. The language of C-symplectic geometry is convenient in context of deformation theory, because it allows us to speak of deformations of a holomorphically symplectic structure without fixing the complex structure beforehand. Using this language, one defines the degenerate twistor deformation as follows.
Theorem 1.2:  Let $M\stackrel{{\scriptstyle\pi}}{{{\>\longrightarrow\>}}}X$ be a Lagrangian fibration on a holomorphically symplectic manifold $(X,\Omega)$, and $\eta\in\Lambda^{2}(X)$ a closed $(2,0)+(1,1)$-form on the base. Then the forms $\Omega_{t}=\Omega+t\pi^{*}\eta$ on $X$ are C-symplectic, and this deformation (called the degenerate twistor deformation) preserves the Lagrangian fibration and the base.
Proof: [BDV, Theorem 3.3].
Using this observation and arguments of linear-algebraic nature, Bogomolov, Deev and Verbitsky proved the following theorem.
Theorem 1.3:  Let $\pi:\;M{\>\longrightarrow\>}X$ be a Lagrangian fibration on a holomorphic symplectic manifold $(M,\Omega)$, and $S:\;X{\>\longrightarrow\>}M$ a smooth ($C^{\infty}$) section. Denote by $\eta\in\Lambda^{2}(X)$ the form $-S^{*}(\Omega)$. Then $\Omega+\pi^{*}\eta$ is C-symplectic, and, moreover, the image of $S$ is holomorphic in the complex structure $I_{\eta}$ induced by the C-symplectic form $\Omega+\pi^{*}\eta$.
Proof: [BDV, Theorem 3.5].
Remark 1.4: 1.2 follows from 1.2 and an observation of N. Hitchin ([Hit, Proposition 1]), who proved that any submanifold $N\subset(M,\Omega)$ of a holomorphically symplectic manifold, which is Lagrangian with respect to $\operatorname{Re}\Omega$ and $\operatorname{Im}\Omega$, is holomorphic Lagrangian. Indeed, the restriction of $\Omega+\pi^{*}\eta$ to $S(X)$ vanishes by construction.
After 1.2 was obtained, it seemed that existence of holomorphic sections of Lagrangian fibrations is easily reachable by topological arguments. After years of painstaking effort, this illusion shattered. In fact, the topological arguments don’t work well even for K3 surfaces. In many years of trying we were unable to construct a smooth section of an elliptic fibration on a K3 surface without using arguments from algebraic geometry. The aim of this paper is to develop this algebro-geometric approach further, in order to solve the topological problem of constructing a smooth section.
1.3 Shafarevich-Tate deformations and the relative Albanese scheme
The Shafarevich-Tate group has its origins in number theory, but in the 1980-ies it appeared in the context of complex algebraic geometry and topology, in the works of R. Friedman and J. Morgan on elliptic surfaces ([FM]). Fridman and Morgan have defined the “analytic Tate-Shafarevich group” as the first cohomology of the sheaf of fiberwise automorphisms of an elliptic surface over its base. Using an explicit computation of this group, they proved that any elliptic surface can be deformed to a surface with a section.
The first attempt to generalize this notion to hyperkähler manifolds appeared in a work of E. Markman [Mar]; Markman produced a family of deformations of a hyperkähler manifold of K3[n]-type with a Lagrangian fibration, with the same fibers and the same base. Markman has explored the Kählerness of the deformation, and shown that in some situations the Kähler property is preserved. His construction was later generalized to all hyperkähler manifolds by A. Abasheva and V. Rogov, ([AR, A]) who also generalized his result about the Kählerness. Also, in [AR, Section 6], the existence of holomorphic sections of a degenerate twistor deformation of a Lagrangian fibration was interpreted as vanishing of a certain cohomology group related to the Shafarevich-Tate group of this Larangian fibration.
The paper [V1] where the degenerate twistor deformations were defined appeared at the same time as Markman’s [Mar]; however, it was unknown what is the precise relation between these two classes of deformations. The Kählerness of the degenerate twistor deformations was finally proven in [SV2].
After the first version of this paper was published, several important works dealing with the subject appeared. In [Sa], G. Saccá defined the relative Albanese variety, a relative group scheme of fiberwise automorphisms of Lagrangian fibration on a hyperkähler manifold without multiple fibers. The Shafarevich-Tate deformations defined by Markman and Abasheva-Rogov naturally appear as cohomology of the sheaf of local sections of this relative group scheme. Another construction of the relative Albanese variety appeared in [Ki] by Y.-J. Kim, who used a multi-dimensional version of the Néron model.
One of the results of G. Saccà [Sa] is especially relevant to the present work. Using the minimal model program, she proves that the relative Albanese variety of a hyperkähler manifold $M$ admits a smooth hyperkähler compactification. This gives a manifold with a Lagrangian fibration admitting a holomorphic section; however, it is not clear whether this new manifold is deformationally equivalent (or even diffeomorphic) to $M$.
After the third version of this paper was published in arxiv.org, J. Kollar produced a version of our main result,[Ko3]. In this paper, he shows that for any Abelian fiber space there exists a Shafarevich-Tate twist which has a birational section. G. Saccà conjectured that any Shafarevich-Tate twist of a hyperkähler manifold $M$ is deformational equivalent to $M$ ([Sa, Remark 6.14]), and proved that it has the same rational cohomology ([Sa, Proposition 6.15]).
1.4 Scheme of the proof
In this subsection, we give a rough outline of the arguments used in this paper.
An integer homology class, generally speaking, cannot be represented by a smooth submanifold. However, for any smooth manifold $M$, a class $z\in H_{r}(M,{\mathbb{Z}})$ can be represented by a submanifold when $r\leqslant 5$. Such results were obtained for different dimensions and classes of embeddings by R. Thom, [T]. The strongest result of this type is the following theorem.
Theorem 1.5:  For any orientable smooth $n$-manifold $V$, all elements of the following integral homology groups can be realized by orientable submanifolds: $H_{n-1}(V,{\mathbb{Z}})$, $H_{n-2}(V,{\mathbb{Z}})$, $H_{i}(V,{\mathbb{Z}})$ for all $i\leqslant 5$.
Proof: [NT, Theorem II.27].
In conjunction with the theory of Néron models, Thom’s result is used to prove that, in the assumptions of 1.1, for any Lagrangian fibration $\pi:\;M\to X$, there exists a smooth 2-dimensional submanifold $S\subset M$ such that the Lagrangian projection $\pi$ projects $S$ to ${\mathbb{C}}P^{1}\subset X$ bijectively; in other words, we obtain a smooth section of $\pi$ over a holomorphic line $C={\mathbb{C}}P^{1}\subset M$ (3.1).
Taking a small neighbourhood of $C$, we can extend this section to a small neighbourhood $U_{C}$, obtaining a map $\varphi:\;U_{C}{\>\longrightarrow\>}\pi^{-1}(U_{C})$. Applying 1.2, we immediately obtain a degenerate twistor deformation of the Lagrangian projection $\pi^{-1}(U_{C})\stackrel{{\scriptstyle\pi}}{{{\>\longrightarrow\>}}}U_{C}$ which makes this section holomorphic.
If we could extend this degenerate twistor deformation to $M$, we would have been able to obtain a holomorphic section of $\pi$ over a line $C\subset X$. However, to extend the twistor deformation from $U_{C}$ to $X={\mathbb{C}}P^{n}$, we need to extend the closed $(2,0)+(1,1)$-form $\eta:=-\varphi^{*}(\Omega)$ to a closed $(2,0)+(1,1)$-form on ${\mathbb{C}}P^{n}$. Such an extension, generally speaking, has an obstruction, which is an element in $H^{1}({\cal O}_{C_{U}})$ (Subsection 5.1), and this group is infinite dimensional, as follows from a simple computation.
In Subsection 5.3, we compute this obstruction class, realizing it by a current of special type, which we call a Dolbeault current, and prove that it vanishes (5.3).
This allows us to construct a degenerate twistor deformation $(M^{\prime},\Omega^{\prime})$ and a holomorphic section of $\pi:\;M^{\prime}{\>\longrightarrow\>}X$ over a neighbourhood of a line $C\subset X={\mathbb{C}}P^{1}$. We notice that $C$ is an ample curve, that is, a curve with an ample normal bundle. For an ample curve $C\subset X$, any holomorphic map $\varphi:\;U_{C}{\>\longrightarrow\>}Z$ from its neighbourhood $U_{C}$ to a complex variety $Z$ can be extended to a meromorphic map $\widetilde{\varphi}:\;X{\>\longrightarrow\>}Z$ (4.3), and this finishes the proof.
1.5 Multiple fibers and primitivity
The assumption of primitivity of the general fiber of Lagrangian projection is clearly necessary for existence of bimeromorphic section. In early versions of this paper, we tried to treat it by using the notion of multiple fibers in codimension 1: conjecturally, no Lagrangian fibration on a hyperkähler manifold may have multiple fibers in codimension 1. Later, we became aware of the results of I. Hellmann [Hel], who constructed a Lagrangian family with multiple fibers (in codimension bigger than 1). After analyzing this example, it became obvious that we use the primitivity of the homology class of the general fiber of the Lagrangian projection. In the current version of this paper, we made this depencency more explicit.
We start by explaining the terminology, which in some of published literature might depend on conventions.
Definition 1.6: Let $\pi:\;M{\>\longrightarrow\>}X$ be a proper holomorphic map, $x\in X$ a point, $F_{x}:=\pi^{-1}(x)$ and $F_{i}$ its irreducible component, with (scheme-theoretical) multiplicity $\mu_{i}$. Denote the greatest common divisor of $\mu_{i}$ by $\mu$. A fiber is multiple if $\mu>1$. A fiber $F_{x}$ is reduced if $\mu_{i}=1$ for all $i$. A fiber $F_{x}$ is has a reduced component if $\mu_{i}=1$ for at least one $i$.
The discriminant of $\pi$ is the set of critical values of $\pi$. By a theorem of Hwang and Oguiso ([HO]), the discriminant is a divisor in $X$. We say that the fibers of $\pi$ have a reduced component in codimension 1 if the fiber $\pi^{-1}(s)$ is not multiple outside of a codimension 2 subset of $X$. It was conjectured that in codimension 1, the fibers of any Lagrangian fibration are actually reduced. In an earlier version of this paper, we argued that the assumption of having a reduced component for the fibers in codimension 1 is sufficient to have a deformation admitting a bimeromorphic section.
It is clear that when a bimeromorphic section exists, the cohomology class of a general fiber should be primitive, and this rules out the existence of multiple fibers. We still hope that the bimeromorphic section exists when the fibers of $\pi$ have a reduced component in codimension 1.
Let $M\to{\mathbb{C}}P^{n}$ be a Lagrangian fibration on a hyperkähler manifold, and $C\subset{\mathbb{C}}P^{n}$ a general line. We use the primitivity of a fiber in Section 3, when we argue for the existence of a 2-dimensional submanifold $N$ in $M_{1}:=\pi^{-1}(C)$ such that $N\cap\pi^{-1}(pt)=1$, where $\cap$ means the intersection pairing in $H_{*}(M_{1})$. By R. Thom’s theorem (1.4), the existence of such $N$ is equivalent to primitivity of $\pi^{*}([H])$ in $H^{2}(M_{1})$, where $[H]$ denotes the fundamental class of the hyperplane section.
Another place where the primitivity of the general fiber is used is Subsection 5.3. In the first step of the proof of 5.3, we invoke the same argument as used in Section 3 to produce the section over a curve. Here we apply it to extend this section to a piecewise smooth multisection of $\pi:\;M\to{\mathbb{C}}P^{n}$, in order to prove that the Dolbeault current associated with the section of $\pi$ in a neighbourhood of a line vanishes.
In [KV], Kamenova and Verbitsky have proven that $\pi^{*}([H])$ is primitive in $H^{*}(M)$, assuming that $\pi:\;M\to{\mathbb{C}}P^{n}$ has no multiple fibers in codimension 1. However, it does not imply that $\pi^{*}([H])$ is primitive in $H^{*}(M_{1})$. Indeed, in [Hel] I. Hellmann has shown that the Mukai system $M$ of rank two and genus two on a K3 surface admits a Lagrangian fibration $\pi:\;M\to X$ with a (very special) fiber of multiplicity 2. Therefore, $\pi$ has no bimeromorphic sections. If $\pi^{*}([H])$ were primitive in $H^{*}(M_{1})$, 1.1 would imply that $\pi$ has no multiple fibers.
However, existence of a section of $\pi:\;M\to{\mathbb{C}}P^{n}$ over a line is not sufficient. Indeed, in Hellmann’s example, such a section exists (it exists for all Beauville-Mukai Lagragian projections). For the benefit of the reader, we explain this construction now. [^1] This argument is due to Giulia Saccá.
Let $M$ be a K3 surface. The Beauville-Mukai manifold is an irreducible component of the moduli space of coherent sheaves obtained as deformations of a line bundle of degree $d$ on a curve of genus $g$ in a K3 surface.
Fix a Beauville-Mukai system of all curves of genus $g$ in $M$, equipped with a line bundle of degree $d$. Appropriately compactified, this gives a hyperkähler manifold $X$ and a Lagrangian fibration $\pi:\;X\to Y$, where $Y={\mathbb{C}}P^{n}$ is the linear system of all curves of genus $g$ on $M$, and $\pi$ is the forgetful map taking the sheaf to its support variety with appropriate multiplicity.
Consider a rational pencil $C$ of curves of genus $g$ passing through a fixed $z\in M$. This is a rational curve on $Y$, and the Beauville-Mukai Lagrangian family has a section $\sigma$ over $C$, which takes a genus $g$ curve $S\subset M$ and picks the line bundle $L={\cal O}(d[z])$; where $d[z]$ denotes a divisor on $S$, obtained by taking $z$ $d$ times,
However, it is impossible to extend $\sigma:\;C\to X$ to a holomorphic section over the tubular neighbourhood of $C$. A smooth section, which necessarily exists, would violate 5.3, that is, the corresponding Dolbeault current is non-zero; in particular, it is impossible to extend the section $\sigma$ to a rational section of $\pi:\;X\to Y$, for any degenerate twistor deformation of $X$.
For Hellmann’s example, the homology class of the general fiber of the Lagrangian fibration is manifestly non-primitive, hence a bimeromorphic section cannot exist. We are grateful to C. Voisin, G. Saccá and A. Abasheva for this observation.
In [Ko3], J. Kollár proved a general result which allows one to construct rational sections of abelian fibrations. His result is valid for any abelian fibration with numerically trivial relative canonical bundle (this case covers all hyperkähler and Calabi-Yau manifolds). Instead of degenerate twistor transform (which makes no sense outside of holomorphic symplectic geometry), Kollár works with the Tate-Shafarevich twist of an Abelian fibration $\pi:\;M\to X$; it is a fibration with the same base and the same fibers, obtained from a 1-cocycle on $X$ with coefficients in fiberwise automorphisms of $M$. Let $F$ be a general fiber of $\pi$ and $d:=\dim M-\dim X$. Kollár has shown that, given a homology class $\eta\in H_{2d}(M,{\mathbb{Z}})$ such that $\eta\cap F=1$, there exists a Tate-Shafarevich twist $\pi_{\eta}:\;M_{\eta}\to X$ and a rational section of this fibration.
The existence of such $\eta$ implies non-existence of multiple fibers. However, unlike the degenerate twistor transform, the Tate-Shafarevich twist of $M$ does not need to be deformationally equivalent (or even diffeomorphic) to $M$.
It is worth comparing our main result to results of [Ko3]. In some of the earlier versions, we assumed that the Lagrangian fibration has reduced fibers in codimension 1. In the current version, we drop this assumption, following the idea of Kollár (3.2). However, our final statement is different from the one in [Ko3], because we use the degenerate twistor deformation, which does not change the deformational type of a manifold, and Kollár uses the Tate-Shafarevich twist, for which we do not know whether it preserves the deformation class. On the other hand, [Ko3] applies to all fibrations with abelian fibers, and our result works only for Lagrangian fibrations on hyperkähler manifolds.
2 Preliminaries on holomorphic symplectic geometry
2.1 C-symplectic structures
In this subsection, we introduce the formalism of C-symplectic structures, used further on to explore the degenerate twistor deformations. We follow [BDV] and [SV1].
Definition 2.1: Let $V$ be a real vector space. A complex structure operator on $V$ is an element $I\in\operatorname{Hom}(V,V)$, satisfying $I^{2}=-\operatorname{Id}_{V}$.
Claim 2.2: The eigenvalues of $I$ are ±$\pm\sqrt{-1}\>$. Moreover, $I$ is diagonalizable over ${\mathbb{C}}$.
Definition 2.3: Let $V$ be a vector space, and $I\in\operatorname{End}(V)$ be a complex structure operator. The eigenvalue decomposition $V\otimes_{\mathbb{R}}{\mathbb{C}}=V^{1,0}\oplus V^{0,1}$ is called the Hodge decomposition; here $I{\left|{}_{{\phantom{|}\!\!}_{V^{1,0}}}\right.}=\sqrt{-1}\>\operatorname{Id}$, and $I{\left|{}_{{\phantom{|}\!\!}_{V^{0,1}}}\right.}=-\sqrt{-1}\>\operatorname{Id}$.
Remark 2.4: One can reconstruct $I$ from the space $V^{1,0}\subset V\otimes_{\mathbb{R}}{\mathbb{C}}$. Indeed, take $V^{0,1}=\overline{V^{1,0}}$, and let $I$ act on $V^{0,1}$ as $\sqrt{-1}\>\operatorname{Id}$, and on $V^{0,1}$ as $-\sqrt{-1}\>\operatorname{Id}$. Since the operator $I\in\operatorname{End}(V\otimes_{\mathbb{R}}{\mathbb{C}})$ commutes with complex conjugation, it is real, that is, it preserves $V\subset V\otimes_{\mathbb{R}}{\mathbb{C}}$.
Definition 2.5: An almost complex structure on a real $2n$-manifold $M$ is an operator $I\in\operatorname{End}(TM)$ satisfying $I^{2}=-\operatorname{Id}_{TM}$, or equivalently, an $n$-dimensional subbundle $T^{1,0}M\subset TM\otimes_{\mathbb{R}}{\mathbb{C}}$ such that $T^{1,0}M\cap\overline{T^{1,0}M}=0$. The almost complex structure is called integrable (and $M$ is a complex manifold) if $T^{1,0}M$ satisfies $[T^{1,0}M,T^{1,0}M]\subset T^{1,0}M$.
Remark 2.6: Integrability implies existence of complex coordinates (by the classical theorem of Newlander-Nirenberg).
Definition 2.7: Let $(M,I)$ be a complex manifold, and let $\Omega\in\Lambda^{2}(M,{\mathbb{C}})$ be a differential form. We say that $\Omega$ is non-degenerate if $\ker\Omega\cap T_{\mathbb{R}}M=0$. We say that it is holomorphically symplectic if it is non-degenerate, $d\Omega=0$, and $\Omega(IX,Y)=\sqrt{-1}\>\Omega(X,Y)$.
Remark 2.8: The equation $\Omega(IX,Y)=\sqrt{-1}\>\Omega(X,Y)$ means that $\Omega$ is complex linear with respect to the complex structure on $T_{\mathbb{R}}M$ induced by $I$.
Remark 2.9: Consider the Hodge decomposition $T_{\mathbb{C}}M=T^{1,0}M\oplus T^{0,1}M$ (the decomposition according to $I$’s eigenvalues). Since $\Omega(IX,Y)=\sqrt{-1}\>\Omega(X,Y)$ and $I(Z)=-\sqrt{-1}\>Z$ for any $Z\in T^{0,1}(M)$, we have $\ker(\Omega)\supset T^{0,1}(M)$. Since $\ker\Omega\cap T_{\mathbb{R}}M=0$, the real dimension of the kernel is at most $\dim_{\mathbb{R}}M$, giving $\dim_{\mathbb{R}}\ker\Omega=\dim M$. Therefore, $\ker(\Omega)=T^{0,1}M$.
Corollary 2.10: Let $\Omega$ be a holomorphically symplectic form on a complex manifold $(M,I)$. Then $I$ is determined by $\Omega$ uniquely.
Claim 2.11: Let $M$ be a smooth $2n$-dimensional manifold. Then there is a bijective correspondence between the set of almost complex structures and the set of sub-bundles $T^{0,1}M\subset TM\otimes_{\mathbb{R}}{\mathbb{C}}$ satisfying $\dim_{\mathbb{C}}T^{0,1}M=n$ and $T^{0,1}M\cap TM=0$ (the last condition means that there are no real vectors in $T^{1,0}M$, that is, that $T^{0,1}M\cap T^{1,0}M=0$).
Proof: Set $I{\left|{}_{{\phantom{|}\!\!}_{T^{1,0}M}}\right.}=\sqrt{-1}\>$ and $I{\left|{}_{{\phantom{|}\!\!}_{T^{0,1}M}}\right.}=-\sqrt{-1}\>$.
Definition 2.12: ([BDV]) Let $M$ be a smooth $4n$-dimensional manifold. A closed complex-valued form $\Omega$ on $M$ is called C-symplectic if $\Omega^{n+1}=0$ and $\Omega^{n}\wedge\overline{\Omega}^{n}$ is a non-degenerate volume form.
Theorem 2.13: Let $\Omega\in\Lambda^{2}(M,{\mathbb{C}})$ be a C-symplectic form, and let $T^{0,1}_{\Omega}(M)$ be equal to $\ker\Omega$, where
$\ker\Omega:=\{v\in TM\otimes{\mathbb{C}}\ \ |\ \ i_{v}(\Omega)=0\},$
where $i_{v}$ denotes contraction with $v$. Then $T^{0,1}_{\Omega}(M)\oplus\overline{T^{0,1}_{\Omega}(M)}=TM\otimes_{\mathbb{R}}{\mathbb{C}}$, hence the sub-bundle $T^{0,1}_{\Omega}(M)$ defines an almost complex structure $I_{\Omega}$ on $M$. If, in addition, $\Omega$ is closed, then $I_{\Omega}$ is integrable, and $\Omega$ is holomorphically symplectic on $(M,I_{\Omega})$.
Proof: The rank of $\Omega$ is $2n$ because $\Omega^{n+1}=0$ and $\Omega^{n}\neq 0$ everywhere. Then $\ker\Omega\oplus\overline{\ker\Omega}=T_{\mathbb{C}}M$. The relation $[T^{0,1}_{\Omega}(M),T^{0,1}_{\Omega}(M)]\subset T^{0,1}_{\Omega}(M)$ follows from Cartan’s formula
$d\Omega(X_{1},X_{2},X_{3})=\frac{1}{6}\sum_{\sigma\in\Sigma_{3}}(-1)^{\widetilde{\sigma}}\operatorname{Lie}_{X_{\sigma_{1}}}\Omega(X_{\sigma_{2}},X_{\sigma_{3}})+(-1)^{\widetilde{\sigma}}\Omega([X_{\sigma_{1}},X_{\sigma_{2}}],X_{\sigma_{3}})$
which gives, for all $X,Y\in T^{0,1}M$, and for any $Z\in TM$,
$d\Omega(X,Y,Z)=\Omega([X,Y],Z),$
implying that $[X,Y]\in T^{0,1}M$.
2.2 Holomorphic Lagrangian fibrations
In this subsection, we introduce holomorphic Lagrangian fibrations, and collect their properties for further use.
Definition 2.14: A holomorphic Lagrangian fibration is a holomorphic map $\pi:\;M{\>\longrightarrow\>}X$ with $M$ holomorphically symplectic and the smooth fibers are Lagrangian with respect to the holomorphic symplectic form.
Theorem 2.15: (Arnold-Liouville) Any smooth fiber of a proper Lagrangian fibration is a torus.
Proof: [E].
Theorem 2.16: (Matsushita, [Mat]) Let $M$ be a hyperkähler manifold of maximal holonomy, and let $\pi:\;M{\>\longrightarrow\>}X$ be a surjective holomorphic map, with $0<\dim X<\dim M$. Then $\pi$ is a Lagrangian fibration.
Theorem 2.17: (Hwang, [Hw]) Under the assumptions in Matsushita’s theorem above, $X$ is biholomorphic to ${\mathbb{C}}P^{n}$ when it is smooth.
Conjecture 2.18: Even under weaker assumptions, i.e., when $X$ is a normal variety, then $X$ is biholomorphic to ${\mathbb{C}}P^{n}$.
Remark 2.19: Originally, Matsushita’s and Hwang’s theorems were only stated and proved in the projective setting. For a most general discussion and proof of these theorems, see [KL, Theorem 2.8] in the setting of singular symplectic varieties, with a normal Kähler variety as the base.
2.3 Degenerate twistor deformations
We round up the preliminaries by exploring the basic properties of degenerate twistor deformations.
Proposition 2.20:  Let $Z\subset M$ be a submanifold of a holomorphically symplectic manifold $(M,\Omega)$, $\dim M=2\dim Z$. Assume that $\Omega{\left|{}_{{\phantom{|}\!\!}_{Z}}\right.}=0$. Then $Z$ is a complex submanifold.
Proof: [Hit, Proposition 1] or [BDV, Proposition 2.7].
In [V1] properties of the degenerate twistor deformation (1.2) on compact hyperkähler manifolds were studied using the Hodge theory.
Remark 2.21: Let $\operatorname{Tw}_{\eta}(M)=M\times{\mathbb{C}}$ be equipped with a complex structure $I_{tw}$ constructed as follows. At any $(x,t)\in M\times{\mathbb{C}}$, $I_{tw}$ acts on $T_{(x,t)}\operatorname{Tw}_{\eta}(M)=T_{x}M\oplus T_{t}{\mathbb{C}}$ as $I_{\Omega_{t\eta}}$ on $T_{x}M$ and as the standard complex structure on $T_{t}{\mathbb{C}}$. It is not hard to see that $I_{tw}$ is integrable. Then $\operatorname{Tw}_{\eta}(M)$ is called the degenerate twistor space. For a hyperkähler manifold, it can be obtained as a limit of twistor spaces ([V1]), which explains the term.
Claim 2.22: ([V1])
The map $\pi:\;M{\>\longrightarrow\>}X$ is holomorphic with respect to $I_{\eta}$, and the complex structures induced by $I$ and $I_{\eta}$ on the fibers of $\pi$ coincide.
2.4 Néron models
We introduce Néron models, starting from the traditional, scheme-theoretic definition (see e. g. [BLR, BHPT]).
The notion of Néron model can be defined in bigger generality, but for most practical purposes, people consider the Néron models of abelian fibrations. In this context, “abelian fibration” is understood as a proper holomorphic map $\pi:\;M{\>\longrightarrow\>}X$ equipped with a relatively ample bundle with an abelian variety as a general fiber and a section. If $X$ is quasiprojective, $\pi$ defines an abelian variety over the field of rational functions on $X$, and this section represents its zero.
Note that the the abelian fibration, as defined above, is equipped with a section, hence it has no multiple fibers.
Definition 2.23: Let $R$ be a Dedekind domain, $K$ its fraction field, and $X$ a scheme over $K$. For a scheme ${\cal Y}$ over $R$ we denote the corresponding scheme ${\cal Y}\times_{\operatorname{Spec}R}\operatorname{Spec}K$ over the fraction field by ${\cal Y}_{K}$. A Néron model of $X$ over $R$ is a smooth, separated finite-type $R$-scheme ${\cal X}$ such that ${\cal X}_{K}=X$ and ${\cal X}$ satisfies the Néron mapping property (NMP): for any smooth $R$-scheme ${\cal Y}$ and any $K$-morphism $f_{K}:\;{\cal Y}_{K}\to{\cal X}_{K}$, $f_{K}$ can be uniquely extended to a morphism $f:\;{\cal Y}{\>\longrightarrow\>}{\cal X}$.
Remark 2.24:  Uniqueness of a Néron model is clear. As follows from [BLR, Theorem 1.4/3], for any abelian variety over $K$, the Néron model always exists. Also, the Néron mapping property implies that the Néron model of an algebraic group over $K$ is a group scheme over $R$.
Remark 2.25: The Néron model was defined over a base which is a spectrum of a Dedekind domain. In the complex geometric setting, this amounts to defining a Néron model over the spectrum of a ring of germs of a smooth complex curve (which is a discrete valuation ring, that is, a local Dedekind domain). This is a local version of this notion, which can be globalized as follows. Define the Néron model of an abelian fibration $\pi:\;Y{\>\longrightarrow\>}C$ over a smooth complex curve $C$ as a complex variety ${\cal Y}\stackrel{{\scriptstyle\varphi}}{{{\>\longrightarrow\>}}}C$ bimeromorphic to $Y$ which gives its Néron model if tensored by the ring ${\cal O}_{C,c}$ of germs of any point $c\in C$.
Note, however, that the Lagrangian fibrations we deal with do not, generally speaking, admit sections, hence they do not define abelian fibrations. If we restrict the fibration to a sufficiently small open set in the base, local sections exist, and we can always define its Néron model.
Theorem 2.26:  Let ${\cal Y}\stackrel{{\scriptstyle\varphi}}{{{\>\longrightarrow\>}}}C$ be an abelian fibration over a smooth complex curve $C$. Assume that ${\cal Y}$ is smooth, algebraic, projective over $C$, and the relative canonical bundle of ${\cal Y}/C$ is trivial. Then the Néron model of ${\cal Y}_{K}$ is isomorphic to the smooth locus of $\varphi$, which is an open subset in ${\cal Y}$.
Proof: In an earlier version of this paper, we followed [BHPT], Section 3 and [HN, Proposition 5.1.8], referring to the notion of a Kulikov model. However, János Kollár [Ko2] pointed out that the definitions of the Kulikov model in these papers are different, and, for the proof, [BHPT] refers to a claim in [HN], which does not exist.
In the present version, we use a result of [Ko2]. Recall that a model of ${\cal Y}$ is a variety, fibered over $C$, bimeromorphic to ${\cal Y}$ over $C$, and with generic fiber isomorphic to that of ${\cal Y}$. A minimal model of an abelian fibration is a model which has terminal singularities and numerically trivial relative canonical bundle. Different flavors of this notion are called “Kulikov model” in [BHPT] and [HN, Proposition 5.1.8]. By [Ko2, Corollary 2 (2.2a)], the natural map from the smooth locus of a minimal model to the Néron model of ${\cal Y}$ is an isomorphism. 2.4 is a special case of this theorem.
This result immediately brings the following corollary. It is essentially the only result from this subsection which we are going to use.
Corollary 2.27:  Let ${\cal Y}\stackrel{{\scriptstyle\varphi}}{{{\>\longrightarrow\>}}}C$ be an abelian fibration over a smooth complex curve $C$ admitting local sections. Assume that ${\cal Y}$ is smooth, algebraic, and the relative canonical bundle of ${\cal Y}/C$ is pullback of a line bundle on $C$. Then the smooth locus of $\varphi$ is equipped with a structure of an algebraic group scheme over $C$.
Proof: Locally in $C$, the bundle ${\cal Y}/C$ is trivial, and therefore the smooth locus of $\varphi$ is the Néron model of ${\cal Y}$ (2.4). The Néron model is a group scheme by 2.4.
3 Constructing a smooth section over a curve
3.1 Constructing a smooth section over a curve using Néron models
The main and only aim of this section is the following theorem.
Theorem 3.1:  Let $\pi:\;M\to X$ be a hyperkähler manifold equipped with an Lagrangian fibration, with $X$ smooth (and hence, by Hwang’s theorem, isomorphic to ${\mathbb{C}}P^{n}$), and $C\subset X$ a general line in $X={\mathbb{C}}P^{n}$. Assume that the general fiber of the Lagrangian projection $\pi$ is primitive (that is, not divisible) in integer homology. Assume, moreover, that $\pi:\;\pi^{-1}(C)\to C$ has local sections. Then there exists a smooth section $\sigma:\;C{\>\longrightarrow\>}M$.
Remark 3.2: The existence of local sections of $\pi:\;\pi^{-1}(C)\to C$ is needed because we use a Néron model of the fibration $\pi:\;\pi^{-1}(C)\to C$. In the next subsection we give another version of this theorem (3.2), which does not use Néron models and does not need this assumption.
Proof. Step 1: By Bertini’s theorem, the preimage $M_{1}:=\pi^{-1}(C)$ of $C$ is smooth. Denote by $Z\subset M_{1}$ the preimage of a general point. We are going to show that its fundamental class $[Z]\in H_{2}(M_{1})$ is primitive in the integral cohomology, that is, is not divisible by a non-invertible integer. Let $F:=\pi^{-1}(pt)$ be a general fiber of $\pi:\;M{\>\longrightarrow\>}{\mathbb{C}}P^{n}$, and $[F]\in H_{2n}(M,{\mathbb{Z}})$ its homology class. Since $[F]$ is primitive, there exists a homology class $\alpha\in H_{2n}(M,{\mathbb{Z}})$ such that $[F]\cap\alpha=1$. Consider the class $\beta:=\alpha\cap[M_{1}]\in H_{2}(M_{1},{\mathbb{Z}})$; it can be realized as an intersection of $\alpha$ and $M_{1}$, if the chains defining $\alpha$, are transversal to $M_{1}$. By construction, this class intersects $[F]$ in one point, if counted with multiplicities, which implies that $[F]\cap\beta=1$.
Essentially, this argument boils down to the following: if $\pi$ has a singular (as in “singular chain”) multisection $\alpha$ of total multiplicity 1, then the intersection of $\alpha$ with $M_{1}$ is a singular chain which gives a multisection of the map $\pi:\;M_{1}{\>\longrightarrow\>}C$ of total multiplicity 1.
Step 2: By the classical result of R. Thom (1.4), $\eta_{1}\in H^{2}(M_{1},{\mathbb{Z}})$ can be represented by an embedded 2-submanifold $S\subset M_{1}$. We can assume that $S$ does not meet the singular locus of $\pi:\;M_{1}{\>\longrightarrow\>}C$ by dimensional reasons. Then, using the transversality theorem, we can also assume that $S$ intersects the fibers of $\pi$ in finitely many points. Since $\eta_{1}\cap F=1$, where $F$ is a fiber, these points add up to 1, if counted with multiplicities. Let $T$ be a smooth fiber of $\pi:\;M_{1}{\>\longrightarrow\>}C$; by Arnold-Liouville’s theorem (2.2), $T$ is a torus. Denote the points of $S\cap T$ by $x_{1},...,x_{k+1},y_{1},...,y_{k}$, where $x_{i}$ are points of positive multiplicity, and $y_{i}$ of negative multiplicity. Fix a reference point $e\in T$. This turns $T$ into a compact abelian group $T_{0}$. Consider the point
$z:=e+\sum_{i=1}^{k+1}(x_{i}-e)-\sum_{i=1}^{k}(y_{i}-e)$ (3.1)
We are going to show that $z$ is independent of the choice of $e\in T$. Clearly, after identifying $T$ and $T_{0}$ by taking $e$ to be the origin, we obtain that $z=\sum_{i=1}^{k+1}x_{i}-\sum_{i=1}^{k}y_{i}$. However, (3.1) describes the dependence on the choice of $e$. Indeed, we can consider $T$ as a manifold equipped with a free and transitive action of $T_{0}$, which is independent of $e$; the expressions $(x_{i}-e)$ and $(y_{i}-e)$ belong to $T_{0}$, and $e+\sum_{i=1}^{k+1}(x_{i}-e)-\sum_{i=1}^{k}(y_{i}-e)$ denotes a point in $T$ shifted by these quantities. If we replace $e$ by $e+u$, the sum (3.1) would not change, hence it gives a section of the projection $\pi:\;M_{1}\to C$, defined outside of its discriminant locus.
Step 3: Applying a degenerate twistor deformation, we can always assume that $M_{1}$ is algebraic; indeed, the set of algebraic members of the degenerate twistor deformation is dense in the set of Kähler fibers ([V1]). Let $U\subset M_{1}$ be the smooth locus of $\pi$. Locally over an open subset $W_{C}\subset C$, we can choose a holomorphic section of this fibration, because the fibers of $\pi:\;M_{1}\to C$. Then, by 2.4, the smooth locus of $\pi^{-1}(W_{C})$ is locally identified with its Néron model. The Néron model is a group fibration, with general fiber an abelian variety, and with special fiber an abelian complex group, not necessarily compact. The fiberwise group structure obtained this way is continuous by 2.4. The singular locus $E:=M_{1}\backslash U$ has complex codimension $\geqslant 2$, hence by transversality we can always assume that the manifold $S$ does not intersect $E$. Indeed, the singular locus of $\pi{\left|{}_{{\phantom{|}\!\!}_{M_{1}}}\right.}$ has codimension 1 in the set $\pi^{-1}(C\cap\Delta)$, where $\Delta$ is the discriminant divisor of $\pi$. Since the group structure on the Néron model is continuous, the operation defined in Step 2 can be extended to $S$ considered as a multisection of the Néron model projection to $W_{C}$. This defines a smooth local section of $\pi:\;M_{1}\to C$ over $W_{C}$.
Step 4: It remains to show that the sections that we constructed locally over $W_{C}\subset C$ agree and define a section of $\pi:\;M_{1}\to C$. The possible ambiguity arises from the local choice of the zero section, used in the construction of the Néron model. Clearly, it suffices to check that they agree on the smooth locus of $\pi$. However, outside of the singularities, the fibers of $\pi$ are tori, equipped with an action of the corresponding abelian group, and the operation (3.1) is independent of the choice of the zero.
Remark 3.3: This result was conjectured in [KV, Remark 4.2].
3.2 Constructing a smooth section over a curve (without Néron models)
We give another proof of 3.1 using the idea communicated to us (after the first draft of this paper was published) by János Kollár. This way, we can get rid of the Néron models in 3.1; this also makes it unnecessary to assume that $\pi:\;\pi^{-1}(C)\to C$ has local sections.
Theorem 3.4:  Let $\pi:\;M\to X$ be a hyperkähler manifold equipped with an Lagrangian fibration, with $X$ smooth (and hence, by Hwang’s theorem, isomorphic to ${\mathbb{C}}P^{n}$), and $C\subset X$ a general line in $X={\mathbb{C}}P^{n}$. Assume that the general fiber of the Lagrangian projection $\pi$ is primitive (that is, not divisible) in integer homology. Then there exists a smooth section $\sigma:\;C{\>\longrightarrow\>}M$.
Proof. Step 1: Let $M_{1}:=\pi^{-1}(C)$. Like in the proof of 3.1, Step 2, we can assume that $M_{1}$ contains a smooth submanifold $Z$ such that $[Z]\cap F=1$, where $F=\pi^{-1}(z)$ for a general point $z\in C$. Let $F_{1}$ be a singular fiber of $\pi$. In the first and second step, we show that $Z$ can be chosen in such a way that for some neighbourhood $U\supset F_{1}$, $Z\cap U$ is a disconnected union of holomorphic subvarieties (possibly with an opposite orientation).
If $F_{1}$ is reduced, this observation is clear: we deform $Z$ in such a way that it intersects $F_{1}$ transversally in its smooth points, and notice that every smooth section of the projection $(z_{1},z_{2},...,z_{n})\mapsto z_{1}$ is homotopic to a holomorphic section near a point $(0,z_{2},...,z_{n})$.
Step 2: If $F_{1}$ is not reduced, locally around a smooth point $m$ of its reduction the projection $\pi:\;M_{1}\to C$ can be written as $(z_{1},z_{2},...,z_{n})\mapsto z_{1}^{k}$, where $k$ is the multiplicity of this component of $F_{1}$. Then, as in Step 1, we can deform $Z$ in such a way that it is holomorphic in the coordinates $(z_{1},z_{2},...,z_{n})$ near $m$. In a small neighbourhood of $\pi(m)\in C$, then, $Z$ becomes a ramified cover, isomorphic to the graph of the multivalued function $t\to(\sqrt[k]{t},z_{2},...,z_{n})$.
Step 3: Let $Z_{0}$ be the intersection of $Z$ with the set of all smooth fibers of $\pi:\;M_{1}\to C$. These fibers are complex tori, hence we can apply to $Z_{0}$ the averaging procedure defined in 3.1, Step 2, obtaining a section of $\pi:\;M_{1}\to C$ over the set $C_{0}$ of all regular values of this map. Since near the critical values $Z$ was a holomorphic multivalued function, and the averaging procedure defines a meromorphic map $M_{1}\times_{\mathbb{C}}M_{1}\times_{C}...\times_{C}M_{1}\to M_{1}$, we obtain a section $S$ of $\pi:\;\pi^{-1}(C)\to C$ which is meromorphic near all singular fibers. Since the fibers of $\pi$ are compact, $S$ is holomorphic at each critical value of $\pi:\;\pi^{-1}(C)\to C$ by valuative criterion of properness. Therefore this section is smooth on $C$.
4 Holography principle for ample rational curves
4.1 Barlet spaces: an introduction
Let $Z\subset M$ be a compact subvariety in a complex manifold, and ${\cal Z}$ the set of all complex deformations of $Z$, that is, the set of all complex subvarieties which occur as fibers in a flat holomorphic family of subvarieties containing $Z$. There are two basic ways to equip ${\cal Z}$ with a structure of a complex variety: one could consider the Douady space of subschemes or the Barlet space of cycles in $M$. It won’t make much difference for the present purposes, but we will focus on the Barlet spaces. For an introduction, reference and more details on Barlet spaces, see [Mag, BM].
Equip $M$ with a Riemannian metric; for the sequel, it does not matter which one we chose. Let ${\mathfrak{S}}$ be the set of all closed subsets of $M$. The Hausdorff topology on ${\mathfrak{S}}$ is defined as follows. Fix $\varepsilon>0$ and a closed subset $Z\subset M$. The base of the Hausdorff topology is the set $U_{\varepsilon}(Z)$ of all $Z^{\prime}\in{\mathfrak{S}}$ such that the $\varepsilon$-neighbourhood of $Z$ contains $Z^{\prime}$, and the $\varepsilon$-neighbourhood of $Z^{\prime}$ contains $Z$. It is not hard to see that Hausdorff topology does not depend on the choice of a Riemannian metric.
A complex variety is called equidimensional if all its irreducible components have the same dimension. A subvariety with multiplicities, or a complex analytic cycle is an equidimensional subvariety $Z\subset M$ with “mutiplicities” (positive integer numbers) attached to each irreducible component. We represent complex analytic cycles by a sum $\sum_{i}n_{i}Z_{i}$, where $Z_{i}$ are its irreducible components, and $n_{i}$ the multiplicities. For every complex analytic cycle, its homology class is $\sum_{i}n_{i}[Z_{i}]$, where $Z_{i}\subset M$ is a complex subvariety.
Fix an integer homology class $\eta\in H_{2i}(M,{\mathbb{Z}})$. The Barlet space parametrizes complex analytic cycles in $M$ with a fixed homology class. The Barlet space is equipped with the topology which is induced by the Hausdorff topology on the set of subsets. As shown in e.g. [BM], the Barlet space is equipped with a complex structure, which is compatible with the Hausdorff topology.
Definition 4.1: The Douady space is the moduli of equidimensional dimensional complex subspaces of $M$ (“subspaces” here means that we allow nilpotents in the structure sheaf). The Hilbert-Chow morphism is the forgetful map $F:\;{\cal D}{\>\longrightarrow\>}{\cal B}$ from the Douady space to the Barlet space, with multiplicity of each component of $\varphi(\widetilde{Z})$ given by the rank of the structure sheaf of $Z$ in the general point of this component.
Example 4.2: For $\dim Z=0$, the Douady space is the Hilbert space $M^{[n]}$ of 0-dimensional subschemes, and the Hilbert-Chow morphism is the natural projection $M^{[n]}{\>\longrightarrow\>}\operatorname{Sym}^{n}M$.
For the present purposes, the following theorem is most important. Fix a Hermitian form $\omega$ on $M$. Recall that the volume $\operatorname{Vol}_{\omega}(Z):=\int_{Z}\omega^{\dim_{\mathbb{C}}Z}$ is well defined for any complex subvariety, even singular, finite when $Z$ is compact, and equal to the cohomology pairing $\langle[Z],\omega^{\dim_{\mathbb{C}}Z}\rangle$ when $\omega$ is closed ([Sto]). For each complex analytic cycle $Z=\sum_{i}n_{i}Z_{i}$ of dimension $k$, we define its Hermitian volume as $\operatorname{Vol}_{\omega}(Z)=\sum_{i}n_{i}\int_{Z_{i}}\omega^{k}$. It is easy to see that the function $\operatorname{Vol}_{\omega}$ is continuous on the Barlet space.
Theorem 4.3:  (E. Bishop’s compactness theorem)
Let $(M,\omega)$ be a compact Hermitian manifold, and $B_{\eta}$ the Barlet space of all cycles homologous to $\eta\in H_{2k}(M)$. Denote by $B_{\eta}(C)$ the subset of $B_{\eta}$ consisting of all cycles $Z$ with $\operatorname{Vol}_{\omega}(Z)\leqslant C$. Then $B_{\eta}(C)$ is compact.
Proof: [Mag, BM].
Remark 4.4: This theorem also holds for pseudoholomorphic curves on compact Hermitian almost complex manifold; in this context it is known as Gromov’s compactness theorem ([Gr]).
4.2 Rationally connected varieties
We give a small introduction to rational connectedness, used in the next subsection. For a proper introduction to this subject, see [C2, KMM].
Definition 4.5: A compact complex manifold $M$ is called rationally connected if any generic $x,y\in M$ can be connected by a sequence of intersecting rational curves, and all irreducible components of Barlet space of curves on $M$ are compact.
Remark 4.6: The condition “all irreducible components of Barlet space of curves on $M$ are compact” would follow, for instance, if $M$ is Fujiki class C, that is, bimeromorphic to Kähler ([F]). However, this condition is more general ([V3]).
Definition 4.7: Recall that a smooth rational curve in a complex manifold is called ample if its normal bundle is $\bigoplus_{i}{\cal O}(k_{i})$, with all $k_{i}$ strictly positive.
Claim 4.8:  Let $M$ be a compact Kähler manifold containing an ample rational curve. Then any $N$ points $z_{1},...,z_{N}$ can be connected by an ample rational curve.
Proof: This fact is well known in algebraic geometry (see [Ko1]). However, its proof is valid for all Kähler manifolds.
The converse statement is also true.
Claim 4.9:  A compact Kähler manifold which contains an ample rational curve is rationally connected.
Proof: See e.g. [H, Lemma 1.2]
Clearly, rational connectedness is invariant under bimeromorphic maps. However, it is not that easy to define rational connectedness for singular varieties. For example, consider a rank 2 holomorphic vector bundle $B$ over $Y$, where $Y$ is a compact complex curve of genus $\geqslant 1$. Let $X:={\mathbb{P}}B$ be its projectivization. It is easy to check that any rational curve in $X$ is projected to a point in $Y$, hence $X$ is not rationally connected. Now, assume that $B=L\oplus{\cal O}_{Y}$, where $L^{*}$ is ample. Then the zero section of the projection $X\to Y$ can be blown down, resulting in a variety $X_{1}$ which can be also obtained as a compactification of the affine cone $C(Y)$. The rational curves in $X$ now all intersect in the singular point of the origin of the cone, but the resolution of $X_{1}$ is not rationally connected.
In the singular setting, definition of rational connectedness can be given as follows. A variety is rationally connected if its resolution of singularities is rationally connected. This is equivalent to the following definition.
Definition 4.10:  Let $M$ be an irreducible compact complex variety $M$ bimeromorphic to a compact Kähler manifold. Then $M$ is called rationally (chain) connected if for any proper subvariety $E\subset M$, any two points in $M$ can be connected by a sequence of intersecting rational curves which intersect in points of $M\backslash E$.
However, for our purposes, an easier criterion is sufficient.
Claim 4.11:  Let $M$ be a compact complex variety, bimeromorphic to a compact Kähler manifold, such that two generic points $x,y\in M$ can be connected by a smooth rational curve $S_{x,y}$. Then $M$ is rationally connected.
Proof: Let $M_{1}\stackrel{{\scriptstyle\pi}}{{{\>\longrightarrow\>}}}M$ be the Kähler resolution of singularities. For $x,y\in M_{1}$ general, the curve $S_{\pi(x),\pi(y)}$ does not belong to the exceptional locus of $\pi$, hence it can be lifted to $M_{1}$. Therefore, $M_{1}$ is rationally connected.
4.3 Holography principle with Kähler target space
To proceed further, we need to generalize the result of [V2], extending the “holography principle” to a Kähler target space. This is the main result of the present section.
Theorem 4.12:  Let $S\subset M$ be an ample rational curve in a $M={\mathbb{C}}P^{n}$, and $U\supset S$ a connected neighbourhood of $S$. Consider an open, connected neighbourhood $U_{S}\subset M$ of $S$. Then any holomorphic map $\varphi:\;U_{S}{\>\longrightarrow\>}X$ from $U_{S}$ to a Kähler manifold $X$ can be extended to a meromorphic map $M\dashrightarrow X$.
Remark 4.13: For 4.3, Kählerness of $X$ is sufficient, but not necessary. As seen from the proof, for 4.3 to hold it suffices to establish that all components of the Barlet spaces of curves on $X$ and $M$ are compact. This holds true if $X$ admits a Hermitian form $\omega$ which is pluriclosed (i.e., it satisfies $dd^{c}\omega=0$), or such that $-dd^{c}\omega$ is a positive form ([V3]).
Remark 4.14: It is not difficult to find a counterexample to 4.3 when $X$ is not simply connected and non-Kähler. For instance, let $X=\operatorname{Tw}(Y)$ be the twistor space for a 2-dimensional complex torus $Y$ (see e.g. [V3]). This is an example of a complex manifold with Barlet components of rational curves non-compact (indeed, in [V3] it is shown that all of these components are Stein). The universal cover of $\operatorname{Tw}(Y)$ is the twistor space for ${\mathbb{C}}^{2}={\mathbb{H}}$, which can be identified with the total space of ${\cal O}(1)^{2}$ ([JV]). Since $\operatorname{Tw}(S^{4})$ is ${\mathbb{C}}P^{3}$ and ${\mathbb{C}}^{2}$ is conformally equivalent to $S^{4}$ without a point, the space $\operatorname{Tw}({\mathbb{C}}^{2})$ is biholomorphic to ${\mathbb{C}}P^{3}\backslash{\mathbb{C}}P^{1}$ (see [JV] for another argument leading to this conclusion). Then an open neighbourhood of ${\mathbb{C}}P^{1}\subset{\mathbb{C}}P^{3}$ can be embedded into $\operatorname{Tw}(Y)$. However, if this embedding can be extended to the whole ${\mathbb{C}}P^{3}$, the space $\operatorname{Tw}(Y)$ would be Moishezon, being an image of a proper holomorphic map from a Moishezon space (see e.g. [F]). This is impossible, because the algebraic dimension of any twistor space of a compact hyperkähler manifold is one ([V2]).
We expect that a similar counterexample would exist when $X$ is non-Kähler but simply connected.
When $X$ is Moishezon, 4.3 is implied by the following theorem (see also [V2, Theorem 3.4]; if we deal with formal schemes instead of complex manifolds, this result was established in [Har] and in [HM, Hir]). The following result requires the manifold $M$ to be homogeneous and rational; for non-homogeneous case, see [V2, Theorem 3.4].
Let $Z\subset M=G/H$ be a subvariety in a complex homogeneous manifold, where $G$ is an algebraic Lie group, $p\in Z$ a point, and $G_{p,Z}$ the set of all elements in $G$ which carry $p$ to a point in $Z$. Denote by $G_{Z}$ the group generated by the set $G_{p,Z}$; clearly, $G_{Z}$ is an algebraic subgroup of $G$, independent from the choice of $p\in Z$. We say that $Z$ generates M if $G_{Z}=G$.
Theorem 4.15:  Let $M$ be a rational homogeneous algebraic variety, $Z$ a complete algebraic subvariety of positive dimension in $M$ which generates $M$, and let $U$ be a connected neighborhood of $Z\subset M$. Then any meromorphic map from $U$ to an algebraic variety $X$ is a restriction of a rational map from $M$ to $X$.
Proof: This is [Chow, Theorem 2, Theorem 3]. Note that “algebraic map” in [Chow, Theorem 2] refers to multi-valued maps given by an algebraic correspondence, which is shown to be rational in [Chow, Theorem 3].
Proof of 4.3: The main idea of the proof of 4.3 is the following observation. Consider an irreducible complex subvariety $B_{0}$ in the Barlet space of curves on $X$ containing $n\varphi(S^{\prime})$, where $n$ is a positive integer (representing multiplicity) and $S^{\prime}\subset U_{S}$ a general deformation of $S$ in $U_{S}$. The corresponding Barlet locus $L_{0}$ is defined as the union of all curves $C\in B_{0}$. Assume that $L_{0}$ is rationally connected. Then $L_{0}$ is Moishezon ([C1]) and its normalization is simply connected ([Deb, Corollary 2.32]; see also [C2, KMM]), hence 4.3 can be applied, showing that the map $\varphi:\;U_{S}{\>\longrightarrow\>}L_{0}$ can be extended to a meromorphic map $M\dashrightarrow L_{0}\subset X$.
Step 1: We take $n=3$. Let $B_{0}$ be the set of all deformations of $3\varphi(S)$ (that is, $\varphi(S)$ with multiplicity 3) which intersect $\varphi(S)$ and intersect $\varphi(U_{S})$ in an open set. The first condition is closed. To get a hold on the second condition, denote by $s$ the intersection of a curve $S_{1}\in B_{0}$ and $\varphi(S)$. Clearly, $S_{1}\cap\varphi(U_{S})$ is open if and only if $S_{1}$ is tangent of infinite order [^1] That is, its $k$-jet in $s$ belongs to the set of $k$-jets of curves in $U_{S}$, for all $k$. to $U_{S}$ in $s$. This condition is closed, because it is obtained by taking an intersection of closed subvarieties enumerated by the order of tangency $k=1,2,3,\dots$. Therefore, $B_{0}$ is closed in the Barlet component $B$ of all deformations of $3\varphi(S)$. From Bishop’s compactness theorem it follows that $B_{0}$ is compact. From Remmert’s proper map theorem, we obtain that the corresponding Barlet locus $L_{0}\subset X$ is complex analytic. To finish the proof, it remains to show that $L_{0}$ is rationally connected. We will prove that two generic points of $L_{0}$ are connected by a smooth curve which belongs to $B_{0}$; by 4.2, this implies the rational connectedness of $L_{0}$.
Step 2: Since $S$ is ample, its normal bundle $NS$ is a direct sum of ${\cal O}(i)$ with $i\geqslant 1$. Therefore, for any two distinct points $x,y\in S$, there exists neighbourhoods $U_{x}\ni x,U_{y}\ni y$ in $U_{S}$ such that any two points $a\in U_{x}$, $b\in U_{y}$ can be connected by a deformation of $S$ which belongs to $U_{S}$. When $NS={\cal O}(1)^{n}$, this curve is unique, hence it cannot intersect $S$ when $a,b$ are generic.
A union of two deformations $S_{1}\ni a,S_{2}\ni b$ of $S$ which intersect can be deformed to a smooth deformation $S_{3}$ of $2S$ by the standard technique ([Ko1]). Moreover, we can perform this deformational smoothing in such a way that $S_{3}$ contains $a$ and $b$. This is the idea used to prove 4.2.
Start from $a,b\in U_{S}$ generic and sufficiently close to $S$, and let $S_{1}$ connect $a$ to $x$, $S_{2}$ connect $b$ to $x$, and $S_{3}$ connect $z\in S$ to $t\in S_{1}$. Smoothing the union $S_{1}\cap S_{2}\cap S_{3}$ as above, we obtain a smooth curve in $U_{S}$ passing through $a,b$ and $z\in S$. By construction, this curve belongs to the Barlet space of deformations of $3S$.
Step 3: Let $L_{0}\subset X$ be the subvariety constructed in Step 1. By construction, $\varphi(U_{S})$ is an open neighbourhood of $S\subset L_{0}$. Therefore, to prove that $L_{0}$ is rationally connected, it would suffice to find a smooth rational curve in $L_{0}$ connecting two general points in $\varphi(U_{S})$. This was done in Step 2. As shown in Step 1, this implies that $L_{0}$ is rationally connected, and finishes the proof of 4.3.
5 Meromorphic sections via Dolbeault currents
5.1 The Dolbeault class of a section
Definition 5.1: Let $\pi:\;M{\>\longrightarrow\>}X$ be a Lagrangian fibration, and $\sigma:\;X{\>\longrightarrow\>}M$ a section. By [BDV, Proposition 2.1] (see also 5.2 below), $\sigma^{*}\Omega$ is a $(2,0)+(1,1)$ closed form on $X$. Assume that $H^{2}(X)$ is generated by a closed $(1,1)$-form $\omega$. Then $\sigma^{*}\Omega-\lambda\omega$ is exact for some $\lambda\in{\mathbb{C}}$. Let $\alpha$ be a 1-form which satisfies $d\alpha=\sigma^{*}\Omega-\lambda\omega$. Then ¯$\overline{\partial}\alpha^{0,1}=0$. The Dolbeaut class of $\sigma$ is the class of $\alpha^{0,1}$ in $H^{1}(\Lambda^{0,*}X,\overline{\partial})=H^{1}(X,{\cal O}_{X})$.
Proposition 5.2:  Let $\pi:\;M{\>\longrightarrow\>}{\mathbb{C}}P^{n}$ be a Lagrangian fibration, $C\subset{\mathbb{C}}P^{n}$ a smooth curve, $C_{U}$ its neighbourhood and $\sigma_{U}:\;C_{U}{\>\longrightarrow\>}M$ a section. Assume that the Dolbeault class $[\alpha^{0,1}]\in H^{1}(C_{U},{\cal O}_{C_{U}})$ of $\sigma_{U}$ vanishes. Then $\sigma^{*}_{U}\Omega$ restricted to a smaller neighbourhood of $C$ can be extended to a closed $(2,0)+(1,1)$-form on ${\mathbb{C}}P^{n}$.
Proof: Let $\alpha$ be a 1-form which satisfies $d\alpha=\sigma^{*}\Omega-\lambda\omega$, where $\omega$ is the Fubini-Study form. Since the $\overline{\partial}$-class of $\alpha^{0,1}$ vanishes, there exists $f\in C^{\infty}M$ such that $\alpha^{\prime}:=\alpha-\overline{\partial}f$ is of type $(1,0)$. Shrinking $C_{U}$ if necessary, we can extend $f$ to a function $\widetilde{f}$ on ${\mathbb{C}}P^{n}$ and $\alpha^{\prime}$ to a $(1,0)$-form $\widetilde{\alpha}^{\prime}$ on ${\mathbb{C}}P^{n}$. Set $\widetilde{\alpha}:=\widetilde{\alpha}^{\prime}+\overline{\partial}\widetilde{f}$. Then $d\widetilde{\alpha}+\lambda\omega$ is a closed $(2,0)+(1,1)$-form on ${\mathbb{C}}P^{n}$ extending $\sigma^{*}\Omega$.
Definition 5.3: Let $U\subset{\mathbb{C}}P^{n}$ be an open subset such that $H^{2}(U)$ is generated by the Fubini-Study form. Let $\rho$ be closed $(2,0)+(1,1)$-form on $U$, and $\rho=d\alpha+\lambda\omega$. Then the $(0,1)$-Dolbeault class of $\rho$ is $[\alpha^{0,1}]\in H^{0,1}(M)$.
Remark 5.4: By definition, the Dolbeault class of $\sigma_{U}$ is equal to the (0,1)-Dolbeault class of $\sigma_{U}^{*}\Omega$.
In the conclusion of this section (Subsection 5.3), we prove the following theorem:
Theorem 5.5: Let $\pi:\;M{\>\longrightarrow\>}{\mathbb{C}}P^{n}$ be a Lagrangian fibration, $C\subset{\mathbb{C}}P^{n}$ a general line, let $C_{U}$ be a neighbourhood of $C$, and $\sigma_{U}:\;C_{U}{\>\longrightarrow\>}M$ a section. Then the Dolbeault class of $\sigma_{U}$ vanishes.
Proof: 5.3.
This theorem completes the last step of the proof of 1.1, the main result of this paper. Before we prove this theorem, we need to make a digression to the land of currents.
5.2 Currents on complex manifolds
For a rich and detailed introduction to currents and their use in algebraic geometry, see [Dem]. We give a few preliminary definitions and cite basic results to facilitate the proof of vanishing of the Dolbeault class of a section, introduced earlier in Subsection 5.1. For all missing proofs, please see [Dem].
Definition 5.6: Let $F$ be a Hermitian bundle with connection $\nabla$, on a Riemannian manifold $M$ with Levi-Civita connection, and
$\|f\|_{C^{k}}:=\sup_{x\in M}\left(|f|+|\nabla f|+...+|\nabla^{k}f|\right)$
the corresponding $C^{k}$-norm defined on smooth sections with compact support. The $C^{k}$-topology is independent from the choice of a connection and a metric.
Definition 5.7: A generalized function is a functional on the top forms with compact support, which is continuous in one of the $C^{i}$-topologies.
Definition 5.8: A $k$-current is a functional on the $(\dim_{\mathbb{R}}M-k)$-forms with compact support, which is continuous in one of the $C^{i}$-topologies.
Remark 5.9: Currents are forms with coefficients in generalized functions.
Definition 5.10: The space of currents is equipped with the weak topology (a sequence of currents converges if it converges on all forms with compact support). The space of currents with this topology is a Montel space (barrelled, locally convex, all bounded subsets are precompact). Montel spaces are reflexive (the map to its double dual with the strong topology is an isomorphism).
Definition 5.11: On a complex manifold, $(p,q)$-currents are $(p,q)$-forms with coefficients in generalized functions.
Remark 5.12: In the literature, this is sometimes called “$(n-p,n-q)$-currents”; we follow the conventions from [Dem].
Claim 5.13: De Rham differential is continuous on currents, and the Poincaré lemma holds for currents, as well as the Poincaré-Dolbeault-Grothendieck lemma. Hence, the cohomology of currents are the same as cohomology of smooth forms, and the $\overline{\partial}$-cohomology groups are the same as for forms.
Definition 5.14: Let $f:\;X{\>\longrightarrow\>}Y$ be a proper holomorphic map of complex manifolds, $\dim_{\mathbb{C}}X=\dim_{\mathbb{C}}Y+k$, and $\alpha$ a $(p,q)$-current on $X$. Define the pushforward $f_{*}\alpha$ using $\langle f_{*}\alpha,\tau\rangle:=\langle\alpha,f^{*}\tau\rangle$, where $\tau$ is any test-form. Then $f_{*}\alpha$ is a $(p-k,q-k)$-current. One should think of $f_{*}$ as of the fiberwise integration.
Remark 5.15: Clearly, $df_{*}\alpha=f_{*}d\alpha$ for any smooth map, $\partial f_{*}\alpha=f_{*}\partial\alpha$ for a holomorphic map, and so on.
Remark 5.16: Pullbacks of currents are (generally speaking) not well-defined.
Using currents, one can give a new proof of the following result of [BDV].
Proposition 5.17:  Let $(M,I,\Omega)$ be a holomorphically symplectic manifold equipped with a holomorphic Lagrangian fibration $\pi:\;M{\>\longrightarrow\>}X$, and let $\sigma:\;X{\>\longrightarrow\>}M$ be a smooth section. Then the pullback $\sigma^{*}(\Omega)$ is a form of type $(2,0)+(1,1)$ on $X$.
Proof: Let $Z:=\operatorname{im}\sigma$, and let $[Z]$ be its integration current, defined by $\langle[Z],\tau\rangle=\int_{Z}\tau$. Clearly, $\sigma^{*}(\Omega)=\pi_{*}(\Omega\wedge[Z])$. Let $\dim_{\mathbb{C}}M=2n$. To prove that $\pi_{*}(\Omega\wedge[Z])$ is of type $(2,0)+(1,1)$ it suffices to show that $\pi_{*}(\Omega\wedge[Z])\wedge\tau=0$ for any $(n,n-2)$-form $\tau$. This would follow if we prove that $\pi^{*}\tau\wedge\Omega=0$. This assertion is clear, because in coordinates $\Omega=\sum dp_{i}\wedge dq_{i}$, where $dp_{i}$ are pullbacks of the coordinates on $X$, and $\tau$ is proportional to $dp_{1}\wedge dp_{2}\wedge...\wedge dp_{n}$.
5.3 Vanishing of the cohomology class of the Dolbeault current
Theorem 5.18:  Let $\pi:\;M{\>\longrightarrow\>}{\mathbb{C}}P^{n}$ be a Lagrangian fibration, $C\subset{\mathbb{C}}P^{n}$ a general line, $C_{U}$ a neighbourhood of $C$, and $\sigma_{U}:\;C_{U}{\>\longrightarrow\>}M$ a section constructed in 3.1. Then the Dolbeault class of $\sigma_{U}$ vanishes.
Proof. Step 1: Let $\pi:\;M{\>\longrightarrow\>}X$ be a Lagrangian fibration, with $\Omega-\pi^{*}\omega=d\alpha$ exact and $\omega$ a closed $(1,1)$-form on $X$. We generalize the Dolbeault class to any piecewise smooth multiple section $Z\subset M$, which is proper and $k$-sheeted (if we count sheets with orientations) over $X$. This is done as follows: we split $Z$ onto smooth pieces, restrict $\alpha$ to each, transport to $X$ with the appropriate sign and sum together, and then take the $\overline{\partial}$-class of the (0,1)-part of resulting current. This is equivalent to taking the current $\pi_{*}(\alpha\wedge[Z])$, using 5.2 to show that its (0,1)-part is $\overline{\partial}$-closed, and taking its $\overline{\partial}$-class.
Step 2: With this definition, we can extend the notion of a Dolbeault class to any multisection of a Lagrangian fibration $\pi:\;M{\>\longrightarrow\>}{\mathbb{C}}P^{n}$, and, indeed, to any piecewise smooth homology chain $u\in H^{2n}(M,{\mathbb{Z}})$, where $\dim_{\mathbb{C}}M=2n$. Defined this way, the Dolbeault class is a homology invariant: for an exact $[Z]=ds$, the pushforward $\pi_{*}(\Omega\wedge[Z])$ satisfies
$\pi_{*}(\Omega\wedge ds)=d(\pi_{*}(\Omega\wedge s))-\pi_{*}(d\Omega\wedge s)=d(\pi_{*}(\Omega\wedge s)).$
Then the Dolbeault class of $\sigma$ is equal to the $(0,1)$-Dolbeault class of the current $d(\pi_{*}(\Omega\wedge s))$, but the (0,1)-Dolbeault class of an exact current vanishes.
Step 3: As shown in 3.1 (step 1 of the proof), the section $\sigma_{C}:\;C{\>\longrightarrow\>}M$ is obtained as a representative of an integral homology class on $M$ intersected with $\pi^{-1}(C)$, hence its homology class can be extended to $M$. This implies that the Dolbeault class of $\sigma_{U}$ vanishes; indeed, a restriction of any class in $H^{0,1}({\mathbb{C}}P^{n})$ to $C_{U}$ vanishes, because $H^{0,1}({\mathbb{C}}P^{n})=0$.
Combining this theorem with 4.3, we obtain the main result of this paper.
Theorem 5.19:  Let $\pi:\;M{\>\longrightarrow\>}{\mathbb{C}}P^{n}$ be a Lagrangian fibration on a hyperkähler manifold. Assume the general fiber of the Lagrangian projection $\pi$ is primitive (that is, not divisible) in integer homology. Then there exists a genenerate twistor deformation $I_{\eta}$ and a complex subvariety $S\subset(M,I_{\eta})$ such that the map $\pi:\;S{\>\longrightarrow\>}{\mathbb{C}}P^{n}$ is birational.
Proof. Step 1: In [SV2], it was shown that any degenerate twistor deformation of a hyperkähler manifold is Kähler. Let $C\subset{\mathbb{C}}P^{n}$ be a general line in the base, and $\sigma:\;C{\>\longrightarrow\>}M$ a smooth section constructed in 3.1. We extend it to an open neighbourhood $C_{U}$ of $C$, obtaining a map $\sigma_{U}:\;C_{U}{\>\longrightarrow\>}M_{U}$, where $M_{U}:=\pi^{-1}(C_{U})$. We apply 1.2 to construct a degenerate twistor deformation $M^{\prime}_{U}$ of $M_{U}$ such that the section $\sigma_{U}:\;C_{U}{\>\longrightarrow\>}M_{U}^{\prime}$ is holomorphic. If this degenerate twistor deformation can be extended from $M_{U}^{\prime}{\>\longrightarrow\>}C_{U}$ to $M^{\prime}{\>\longrightarrow\>}{\mathbb{C}}P^{n}$, then 4.3, applied to the holomorphic map $C_{U}{\>\longrightarrow\>}M^{\prime}$, gives a rational section of $\pi$.
Step 2: Let $\eta_{U}:=\sigma_{U}^{*}(\Omega)$. By 5.2, $\eta_{U}$ is of type $(2,0)+(1,1)$. To obtain $M^{\prime}$, we have to extend $\eta_{U}$ to a closed form of type $(2,0)+(1,1)$ on ${\mathbb{C}}P^{n}$. As seen in Subsection 5.1, there is a cohomological obstruction to such an extension, which belongs to the group $H^{1}(C_{U},{\cal O})$. [^1] It is possible to see that this group is infinite-dimensional.
Step 3: In 5.3 we proved that the obstruction class in $H^{1}(C_{U},{\cal O})$ (i.e., “the Dolbeault class of the section $\sigma_{U}$”) vanishes. Now, 5.1 implies that the degenerate twistor deformation $M_{U}^{\prime}$ can be extended to $M^{\prime}{\>\longrightarrow\>}{\mathbb{C}}P^{n}$.
Acknowledgments. The second named author thanks IMPA for their hospitality in January 2020 when part of this project began. We are grateful to Claire Voisin and Andrey Soldatenkov who helped us to find the error in the first version of this paper. Many thanks to Jason Starr for references to rational connectedness in relation to rational singularities. We are grateful to J. Kollár and G. Saccà for their insightful comments and email exchange.

### v5

@@
 Then for some $t_{0}\in H^{2}(X,{\mathbb{C}})$, the Lagrangian projection $\pi:\;(M,I_{t_{0}}){\>\longrightarrow\>}X$ admits a meromorphic section.
-Proof: 5.3.
+Proof: 6.3.
 1.2 C-symplectic structures and degenerate twistor deformations
@@
 Theorem 1.3: Let $\pi:\;M{\>\longrightarrow\>}X$ be a Lagrangian fibration on a holomorphic symplectic manifold $(M,\Omega)$, and $S:\;X{\>\longrightarrow\>}M$ a smooth ($C^{\infty}$) section.
-Denote by $\eta\in\Lambda^{2}(X)$ the form $-S^{*}(\Omega)$.
+Denote by $\eta\in\Lambda^{2}(X)$ the form $S^{*}(\Omega)$.
 Then $\Omega+\pi^{*}\eta$ is C-symplectic, and, moreover, the image of $S$ is holomorphic in the complex structure $I_{\eta}$ induced by the C-symplectic form $\Omega+\pi^{*}\eta$.
@@
 A fiber $F_{x}$ is has a reduced component if $\mu_{i}=1$ for at least one $i$.
-The discriminant of $\pi$ is the set of critical values of $\pi$.
-By a theorem of Hwang and Oguiso ([HO]), the discriminant is a divisor in $X$.
+The following result should be well-known, but we did not find it in the literature.
+Lemma 1.7: Let $f:\;X{\>\longrightarrow\>}Y$ be a proper holomorphic map of complex manifolds, and $y\in Y$ any point.
+Then $f$ has a local section in a neighbourhood of $y$ if and only if the fiber $f^{-1}(y)$ has a component of multiplicity 1.
+Proof: Assume that $f^{-1}(y)$ has a reduced component $Z$, and consider a transversal to $Z$ local submanifold $U$ of complimentary dimension intersecting $Z$ in a smooth point.
+Since $U\cap f^{-1}(y)=1$, for any $y^{\prime}\in Y$ sufficiently close to $y$, we also have $U\cap f^{-1}(y^{\prime})=1$.
+The implicit function theorem implies that the map $y^{\prime}\mapsto U\cap f^{-1}(y^{\prime})$ is a holomorphic map, defining a local section of $f$.
+Conversely, assume that $f$ has a local section $s:\;V\to X$ in a neighbourhood $V\ni y$.
+Let $s(V)=U$; it is a smooth local submanifold of $X$.
+The intersection $U\cap f^{-1}(y)$ has multiplicity 1, hence it belongs to only one of the irreducible components $Z_{1}$ of $f^{-1}(y)$.
+However, unless $Z_{1}$ is smooth in the intersection point $U\cap Z_{1}$, the intersection $U\cap Z_{1}$ has multiplicity >1>1, which is impossible.
+Therefore, $Z_{1}$ is a smooth submanifold of $X$ in a neighbourhood of the point $U\cap Z_{1}$, and the multiplicity of this component is 1.
+The discriminant of a proper holomorphic map $\pi$ is the set of critical values of $\pi$.
+By a theorem of Hwang and Oguiso ([HO]), the discriminant of a Lagrangian fibration is a divisor in $X$.
 We say that the fibers of $\pi$ have a reduced component in codimension 1 if the fiber $\pi^{-1}(s)$ is not multiple outside of a codimension 2 subset of $X$.
@@
 Another place where the primitivity of the general fiber is used is Subsection 6.3.
-In the first step of the proof of 5.3, we invoke the same argument as used in Section 3 to produce the section over a curve.
+In the first step of the proof of 6.3, we invoke the same argument as used in Section 4 to produce the section over a curve.
 Here we apply it to extend this section to a piecewise smooth multisection of $\pi:\;M\to{\mathbb{C}}P^{n}$, in order to prove that the Dolbeault current associated with the section of $\pi$ in a neighbourhood of a line vanishes.
-In [KV], Kamenova and Verbitsky have proven that $\pi^{*}([H])$ is primitive in $H^{*}(M)$, assuming that $\pi:\;M\to{\mathbb{C}}P^{n}$ has no multiple fibers in codimension 1.
+In [KV1], Kamenova and Verbitsky have proven that $\pi^{*}([H])$ is primitive in $H^{*}(M)$, assuming that $\pi:\;M\to{\mathbb{C}}P^{n}$ has no multiple fibers in codimension 1.
 However, it does not imply that $\pi^{*}([H])$ is primitive in $H^{*}(M_{1})$.
@@
 However, it is impossible to extend $\sigma:\;C\to X$ to a holomorphic section over the tubular neighbourhood of $C$.
-A smooth section, which necessarily exists, would violate 5.3, that is, the corresponding Dolbeault current is non-zero; in particular, it is impossible to extend the section $\sigma$ to a rational section of $\pi:\;X\to Y$, for any degenerate twistor deformation of $X$.
+A smooth section, which necessarily exists, would violate 6.3, that is, the corresponding Dolbeault current is non-zero; in particular, it is impossible to extend the section $\sigma$ to a rational section of $\pi:\;X\to Y$, for any degenerate twistor deformation of $X$.
 For Hellmann’s example, the homology class of the general fiber of the Lagrangian fibration is manifestly non-primitive, hence a bimeromorphic section cannot exist.
@@
 Definition 2.1: Let $V$ be a real vector space. A complex structure operator on $V$ is an element $I\in\operatorname{Hom}(V,V)$, satisfying $I^{2}=-\operatorname{Id}_{V}$.
-Claim 2.2: The eigenvalues of $I$ are ±$\pm\sqrt{-1}\>$. Moreover, $I$ is diagonalizable over ${\mathbb{C}}$.
+Claim 2.2: The eigenvalues of $I$ are $\pm\sqrt{-1}\>$. Moreover, $I$ is diagonalizable over $\mathbb{C}$}.
 Definition 2.3: Let $V$ be a vector space, and $I\in\operatorname{End}(V)$ be a complex structure operator.
@@
 Remark 2.4: One can reconstruct $I$ from the space $V^{1,0}\subset V\otimes_{\mathbb{R}}{\mathbb{C}}$.
-Indeed, take $V^{0,1}=\overline{V^{1,0}}$, and let $I$ act on $V^{0,1}$ as $\sqrt{-1}\>\operatorname{Id}$, and on $V^{0,1}$ as $-\sqrt{-1}\>\operatorname{Id}$.
+Indeed, take $V^{0,1}=\overline{V^{1,0}}$, and let $I$ act on $V^{0,1}$ as $\sqrt{-1}\>\operatorname{Id}$, and on $V^{0,1}$ as $\sqrt{-1}\>\operatorname{Id}$.
 Since the operator $I\in\operatorname{End}(V\otimes_{\mathbb{R}}{\mathbb{C}})$ commutes with complex conjugation, it is real, that is, it preserves $V\subset V\otimes_{\mathbb{R}}{\mathbb{C}}$.
 Definition 2.5: An almost complex structure on a real $2n$-manifold $M$ is an operator $I\in\operatorname{End}(TM)$ satisfying $I^{2}=-\operatorname{Id}_{TM}$, or equivalently, an $n$-dimensional subbundle $T^{1,0}M\subset TM\otimes_{\mathbb{R}}{\mathbb{C}}$ such that $T^{1,0}M\cap\overline{T^{1,0}M}=0$.
-The almost complex structure is called integrable (and $M$ is a complex manifold) if $T^{1,0}M$ satisfies $[T^{1,0}M,T^{1,0}M]\subset T^{1,0}M$.
+The almost complex structure is called integrable (and $M$ is a complex manifold) if $T^{1,0}M$ satisfies $T^{1,0}M,T^{1,0}M]\subset T^{1,0}M$.
 Remark 2.6: Integrability implies existence of complex coordinates (by the classical theorem of Newlander-Nirenberg).
@@
 Then $\ker\Omega\oplus\overline{\ker\Omega}=T_{\mathbb{C}}M$.
-The relation $[T^{0,1}_{\Omega}(M),T^{0,1}_{\Omega}(M)]\subset T^{0,1}_{\Omega}(M)$ follows from Cartan’s formula
+The relation $T^{0,1}_{\Omega}(M),T^{0,1}_{\Omega}(M)]\subset T^{0,1}_{\Omega}(M)$ follows from Cartan’s formula
 $d\Omega(X_{1},X_{2},X_{3})=\frac{1}{6}\sum_{\sigma\in\Sigma_{3}}(-1)^{\widetilde{\sigma}}\operatorname{Lie}_{X_{\sigma_{1}}}\Omega(X_{\sigma_{2}},X_{\sigma_{3}})+(-1)^{\widetilde{\sigma}}\Omega([X_{\sigma_{1}},X_{\sigma_{2}}],X_{\sigma_{3}})$
@@
 $d\Omega(X,Y,Z)=\Omega([X,Y],Z),$
-implying that $[X,Y]\in T^{0,1}M$.
+implying that $X,Y]\in T^{0,1}M$.
 2.2 Holomorphic Lagrangian fibrations
@@
 The map $\pi:\;M{\>\longrightarrow\>}X$ is holomorphic with respect to $I_{\eta}$, and the complex structures induced by $I$ and $I_{\eta}$ on the fibers of $\pi$ coincide.
-2.4 Néron models
-We introduce Néron models, starting from the traditional, scheme-theoretic definition (see e. g. [BLR, BHPT]).
-The notion of Néron model can be defined in bigger generality, but for most practical purposes, people consider the Néron models of abelian fibrations.
-In this context, “abelian fibration” is understood as a proper holomorphic map $\pi:\;M{\>\longrightarrow\>}X$ equipped with a relatively ample bundle with an abelian variety as a general fiber and a section.
-If $X$ is quasiprojective, $\pi$ defines an abelian variety over the field of rational functions on $X$, and this section represents its zero.
-Note that the the abelian fibration, as defined above, is equipped with a section, hence it has no multiple fibers.
-Definition 2.23: Let $R$ be a Dedekind domain, $K$ its fraction field, and $X$ a scheme over $K$.
-For a scheme ${\cal Y}$ over $R$ we denote the corresponding scheme ${\cal Y}\times_{\operatorname{Spec}R}\operatorname{Spec}K$ over the fraction field by ${\cal Y}_{K}$.
-A Néron model of $X$ over $R$ is a smooth, separated finite-type $R$-scheme ${\cal X}$ such that ${\cal X}_{K}=X$ and ${\cal X}$ satisfies the Néron mapping property (NMP): for any smooth $R$-scheme ${\cal Y}$ and any $K$-morphism $f_{K}:\;{\cal Y}_{K}\to{\cal X}_{K}$, $f_{K}$ can be uniquely extended to a morphism $f:\;{\cal Y}{\>\longrightarrow\>}{\cal X}$.
-Remark 2.24:  Uniqueness of a Néron model is clear.
-As follows from [BLR, Theorem 1.4/3], for any abelian variety over $K$, the Néron model always exists.
-Also, the Néron mapping property implies that the Néron model of an algebraic group over $K$ is a group scheme over $R$.
-Remark 2.25: The Néron model was defined over a base which is a spectrum of a Dedekind domain.
-In the complex geometric setting, this amounts to defining a Néron model over the spectrum of a ring of germs of a smooth complex curve (which is a discrete valuation ring, that is, a local Dedekind domain).
-This is a local version of this notion, which can be globalized as follows.
-Define the Néron model of an abelian fibration $\pi:\;Y{\>\longrightarrow\>}C$ over a smooth complex curve $C$ as a complex variety ${\cal Y}\stackrel{{\scriptstyle\varphi}}{{{\>\longrightarrow\>}}}C$ bimeromorphic to $Y$ which gives its Néron model if tensored by the ring ${\cal O}_{C,c}$ of germs of any point $c\in C$.
-Note, however, that the Lagrangian fibrations we deal with do not, generally speaking, admit sections, hence they do not define abelian fibrations.
-If we restrict the fibration to a sufficiently small open set in the base, local sections exist, and we can always define its Néron model.
-Theorem 2.26:  Let ${\cal Y}\stackrel{{\scriptstyle\varphi}}{{{\>\longrightarrow\>}}}C$ be an abelian fibration over a smooth complex curve $C$.
-Assume that ${\cal Y}$ is smooth, algebraic, projective over $C$, and the relative canonical bundle of ${\cal Y}/C$ is trivial.
-Then the Néron model of ${\cal Y}_{K}$ is isomorphic to the smooth locus of $\varphi$, which is an open subset in ${\cal Y}$.
-Proof: In an earlier version of this paper, we followed [BHPT], Section 3 and [HN, Proposition 5.1.8], referring to the notion of a Kulikov model.
-However, János Kollár [Ko2] pointed out that the definitions of the Kulikov model in these papers are different, and, for the proof, [BHPT] refers to a claim in [HN], which does not exist.
-In the present version, we use a result of [Ko2].
-Recall that a model of ${\cal Y}$ is a variety, fibered over $C$, bimeromorphic to ${\cal Y}$ over $C$, and with generic fiber isomorphic to that of ${\cal Y}$.
-A minimal model of an abelian fibration is a model which has terminal singularities and numerically trivial relative canonical bundle.
-Different flavors of this notion are called “Kulikov model” in [BHPT] and [HN, Proposition 5.1.8].
-By [Ko2, Corollary 2 (2.2a)], the natural map from the smooth locus of a minimal model to the Néron model of ${\cal Y}$ is an isomorphism.
-2.4 is a special case of this theorem.
-This result immediately brings the following corollary. It is essentially the only result from this subsection which we are going to use.
-Corollary 2.27:  Let ${\cal Y}\stackrel{{\scriptstyle\varphi}}{{{\>\longrightarrow\>}}}C$ be an abelian fibration over a smooth complex curve $C$ admitting local sections.
-Assume that ${\cal Y}$ is smooth, algebraic, and the relative canonical bundle of ${\cal Y}/C$ is pullback of a line bundle on $C$.
-Then the smooth locus of $\varphi$ is equipped with a structure of an algebraic group scheme over $C$.
-Proof: Locally in $C$, the bundle ${\cal Y}/C$ is trivial, and therefore the smooth locus of $\varphi$ is the Néron model of ${\cal Y}$ (2.4). The Néron model is a group scheme by 2.4.
+3 Canonical models for Lagrangian fibrations
+3.1 Smooth torsor bundle associated with a Lagrangian fibration
+We relate some results about the Néron models and their generalizations to Lagrangian fibered hyperkähler manifolds.
+We follow [KV2].
+The complex-geometric approach used in [KV2] is based on the notion of Kulikov model, discussed in [BLR, BHPT]; it was formally defined in [Ko2] as “canonical model”.
+The main result which we use appeared as a part of Theorem 1.6 in [Ki]. In [KV2] we gave an independent proof of the following statement (see [Ki, Theorem 1.6]).
+Theorem 3.1: Let $M$ be a holomorphic symplectic manifold, and $\pi:\;M\to X$ a Lagrangian fibration, that is, a proper holomorphic map which has holomorphic Lagrangian fibers and admits a relatively ample line bundle.
+Assume that $X$ is an open subset of a projective manifold, and $M_{i}\subset M$ the set of all points where the differental of $\pi$ has rank $\leqslant\dim X-i$.
+[^1] In [KV2], we show that $\operatorname{codim}\pi(M_{i})\geqslant i$.
+Denote by $X_{0}$ the set $X\backslash\pi(M_{2})$, and let $N\subset M$ be $\pi^{-1}(X_{0})\backslash M_{1}$.
+Then the natural projection $N\to X_{0}$ is equipped with a unique structure of a torsor of a smooth group scheme over the base $X_{0}$.
+Proof: [KV2, Theorem 1.1].
+Let us explain the terminology used in this statement.
+Definition 3.2: Let $f:\;X_{0}\to Y$ be a smooth holomorphic submersion (not necessarily proper), and $s:\;Y\to X_{0}$ a section.
+Consider a map $X_{0}\times_{Y}X_{0}\stackrel{{\scriptstyle\mu}}{{\to}}X_{0}$ commuting with the projection to $Y$.
+We say that $(X_{0},s,\mu)$ is a holomorphic group bundle if the map $\mu$ is submersive and defines a group structure on each fiber $f^{-1}(y)$, for all $y\in Y$, and $s(y)$ is its identity element.
+By definition, a torsor over an abelian group $G$ is a set where $G$ acts freely and transitively.
+However, the torsor structure can be defined without referring to an a priori given group structure as follows.
+Definition 3.3: An abelian torsor is a set $S$ equipped with a map $\mu:\;S\times S\times S\to S$ such that
+for any $e\in S$, the map $x\cdot y:=\mu(x,y,e)$ defines a commutative and associative product on $S$ with $e$ an identity element
+for any fixed $y\in S$, the map $x\mapsto x\cdot y$ is bijective.
+Remark 3.4: From the axioms of the abelian torsor, it is clear that $\cdot$ defines a structure of an abelian group on $S$.
+Remark 3.5: The identity element axiom means that $\mu(x,y,y)=x$ for any $x$ and $y$. Given a group structure on $S$, the torsor operation $\mu$ can be defined as $\mu(x,y,z)=x+y-z$.
+The most important property of abelian torsors is used many times throughout this paper.
+Claim 3.6: Let $S$ be a set equipped with a structure of a torsor of an abelian group $G$, and $a_{1},...,a_{n},b_{1},...,b_{n+1}$ points in $S$.
+Consider $e\in S$, and let ++ denote the abelian group structure on $S$ associated with the unit $e$.
+Then the sum $\sum_{i=1}^{n+1}b_{i}-\sum_{i=1}^{n}a_{i}$ is independent from the choice of $e\in S$.
+Proof: Let us denote the group operation on $S$ by $x+_{e}y:=\mu(x,e,y)$, and fix $u\in S$. Then $x+_{u}y=x+_{e}y-_{e}u$, and
+$b_{1}+_{u}b_{2}+_{u}...+_{u}b_{n+1}=b_{1}+_{e}b_{2}+_{e}...+_{e}b_{n+1}-_{e}(n+1)u.$
+Then
+$b_{1}+_{u}b_{2}+_{u}...+_{u}b_{n+1}-_{u}a_{1}-_{u}a_{2}-_{u}...-_{u}a_{n}=$
+$b_{1}+_{e}b_{2}+_{e}...+_{e}b_{n+1}-_{e}a_{1}-_{e}a_{2}-_{e}...-_{e}a_{n}-_{e}(n+1)u+_{e}nu=$
+$b_{1}+_{e}b_{2}+_{e}...+_{e}b_{n+1}-_{e}a_{1}-_{e}a_{2}-_{e}...-_{e}a_{n}$
+(see also the proof of 4.1, Step 2).
+3.2 Néron models for Lagrangian fibrations
+When a Lagrangian fibration $\pi:\;M\to X$ is equipped with a holomorphic section, the torsor bundle defined in 3.1 is actually a group bundle.
+In [Ki], this group bundle was interpreted as a part of the multi-dimensional Néron model of $M$.
+Throughout this paper, we often refer to this construction as to the Néron model construction for the Lagrangian fibration.
+Here is the relevant statement from [KV2] formulated for Lagrangian fibrations equipped with holomorphic sections.
+Definition 3.7: Let $M_{i}$ be the set of all points $m\in M$ such that the differential $D\pi$ has rank $\leqslant\dim X-i$.
+The set $\pi(M_{1})$ of singular values of a Lagrangian fibration map $\pi:\;M\to X$ is called is discriminant, denoted by $\operatorname{Discr}\subset X$.
+Let $\operatorname{Discr}_{i}:=\pi(M_{i+1})$.
+Clearly, $\operatorname{Discr}=\operatorname{Discr}_{0}\supset\operatorname{Discr}_{1}\supset$....
+By [KV2, Claim 2.9], $\operatorname{codim}_{\operatorname{Discr}}\operatorname{Discr}_{i}\geqslant i$.
+Remark 3.8: Note that $\operatorname{Discr}_{i}$ does not necesarily coincide with the $i$-th singular stratum of $\operatorname{Discr}$. However, it has the same dimension.
+Theorem 3.9: Let $M$ be a hyperkähler manifold, $f:\;M\to X$ a Lagrangian fibration, and $M_{0}\subset M$ its smooth locus.
+Denote by $J(f)$ the manifold $M_{0}\cap f^{-1}(X\backslash\operatorname{Discr}_{1})$.
+Assume that $f$ admits a holomorphic section $s:\;X\to M_{0}$.
+Then there exists a unique group bundle structure on $f:\;J(f)\to X\backslash\operatorname{Discr}_{1}$ with the following properties:
+For each $y\in X\backslash\operatorname{Discr}_{1}$, $s(y)$ is the identity element of the fiber $f^{-1}\cap J(f)$.
+[…]
+Proof: [KV2, Theorem 2.13].
 4 Constructing a smooth section over a curve
@@
 Theorem 4.1: Let $\pi:\;M\to X$ be a hyperkähler manifold equipped with an Lagrangian fibration, with $X$ smooth (and hence, by Hwang’s theorem, isomorphic to ${\mathbb{C}}P^{n}$), and $C\subset X$ a general line in $X={\mathbb{C}}P^{n}$.
-Assume that the general fiber of the Lagrangian projection $\pi$ is primitive (that is, not divisible) in integer homology.
-Assume, moreover, that $\pi:\;\pi^{-1}(C)\to C$ has local sections.
-Then there exists a smooth section $\sigma:\;C{\>\longrightarrow\>}M$.
-Remark 3.2: The existence of local sections of $\pi:\;\pi^{-1}(C)\to C$ is needed because we use a Néron model of the fibration $\pi:\;\pi^{-1}(C)\to C$.
-In the next subsection we give another version of this theorem (3.2), which does not use Néron models and does not need this assumption.
-Proof.
-Step 1: By Bertini’s theorem, the preimage $M_{1}:=\pi^{-1}(C)$ of $C$ is smooth.
+Assume that the general fiber $F$ of the Lagrangian projection $\pi$ is primitive (that is, not divisible) in the integral homology.
+Assume, moreover, that the fibers of $\pi$ are reduced away from a codimension $\geqslant 2$ subset of $X={\mathbb{C}}P^{n}$.
+Then
+there exists a smooth section $\sigma:\;C{\>\longrightarrow\>}M$.
+Moreover, this section can be chosen homologous to the homology class $\alpha\cap\pi^{-1}(C)\in H_{2}(\pi^{-1}(C),{\mathbb{Z}})$, where $\alpha\in H_{2n}(M,{\mathbb{Z}})$ a homology class such that $\alpha\cap[F]=1$, and $F]\in H_{2n}(M,{\mathbb{Z}})$ the homology class of a general fiber of $\pi$.
+Remark 4.2: The existence of local sections of $\pi:\;\pi^{-1}(C)\to C$ is needed in 4.3 (ii); it is also used in the original proof of 4.3 (i) based on the Néron models of the fibration $\pi:\;\pi^{-1}(C)\to C$ (given below).
+In Subsection 4.3 we give another version of 4.3 (i), which does not use Néron models and does not need this assumption.
+We end this subsection with a proof of 4.1 (i). In Subsection 4.2, we explain why 4.1 (ii) is non-trivial and prove it, and in Subsection 4.3, we give another proof of 4.1 (i).
+Proof of 4.1 (i), step 1: By Bertini’s theorem, the preimage $M_{1}:=\pi^{-1}(C)$ of $C$ is smooth.
 Denote by $Z\subset M_{1}$ the preimage of a general point.
-We are going to show that its fundamental class $[Z]\in H_{2}(M_{1})$ is primitive in the integral cohomology, that is, is not divisible by a non-invertible integer.
-Let $F:=\pi^{-1}(pt)$ be a general fiber of $\pi:\;M{\>\longrightarrow\>}{\mathbb{C}}P^{n}$, and $[F]\in H_{2n}(M,{\mathbb{Z}})$ its homology class.
-Since $[F]$ is primitive, there exists a homology class $\alpha\in H_{2n}(M,{\mathbb{Z}})$ such that $[F]\cap\alpha=1$.
+We are going to show that its fundamental class $Z]\in H_{2}(M_{1})$ is primitive in the integral cohomology, that is, is not divisible by a non-invertible integer.
+Let $F:=\pi^{-1}(pt)$ be a general fiber of $\pi:\;M{\>\longrightarrow\>}{\mathbb{C}}P^{n}$, and $F]\in H_{2n}(M,{\mathbb{Z}})$ its homology class.
+Since $[F]$ is primitive, there exists a homology class $\alpha\in H_{2n}(M,{\mathbb{Z}})$ such that $F]\cap\alpha=1$.
 Consider the class $\beta:=\alpha\cap[M_{1}]\in H_{2}(M_{1},{\mathbb{Z}})$; it can be realized as an intersection of $\alpha$ and $M_{1}$, if the chains defining $\alpha$, are transversal to $M_{1}$.
-By construction, this class intersects $[F]$ in one point, if counted with multiplicities, which implies that $[F]\cap\beta=1$.
+By construction, this class intersects $[F]$ in one point, if counted with multiplicities, which implies that $F]\cap\beta=1$.
 Essentially, this argument boils down to the following: if $\pi$ has a singular (as in “singular chain”) multisection $\alpha$ of total multiplicity 1, then the intersection of $\alpha$ with $M_{1}$ is a singular chain which gives a multisection of the map $\pi:\;M_{1}{\>\longrightarrow\>}C$ of total multiplicity 1.
@@
 If we replace $e$ by $e+u$, the sum (4.1) would not change, hence it gives a section of the projection $\pi:\;M_{1}\to C$, defined outside of its discriminant locus.
+For an abstract version of this argument (formulated for any abelian torsor) see 3.1.
 Step 3: Applying a degenerate twistor deformation, we can always assume that $M_{1}$ is algebraic; indeed, the set of algebraic members of the degenerate twistor deformation is dense in the set of Kähler fibers ([V1]).
-Let $U\subset M_{1}$ be the smooth locus of $\pi$.
+Let $M_{2}\subset M_{1}$ be the smooth locus of $\pi$.
 Locally over an open subset $W_{C}\subset C$, we can choose a holomorphic section of this fibration, because the fibers of $\pi:\;M_{1}\to C$.
-Then, by 2.4, the smooth locus of $\pi^{-1}(W_{C})$ is locally identified with its Néron model.
-The Néron model is a group fibration, with general fiber an abelian variety, and with special fiber an abelian complex group, not necessarily compact.
-The fiberwise group structure obtained this way is continuous by 2.4.
-The singular locus $E:=M_{1}\backslash U$ has complex codimension $\geqslant 2$, hence by transversality we can always assume that the manifold $S$ does not intersect $E$.
+Then, by 3.1, the smooth locus $M_{2}$ of $\pi^{-1}(W_{C})$ is identified with its torsor Néron model.
+The Néron model is a torsor of a smooth abelian group fibration, with general fiber a compact complex torus, and with special fiber a torsor of an abelian complex group, not necessarily compact.
+The fiberwise torsor structure obtained this way is holomorphic.
+The singular locus $E:=M_{1}\backslash M_{2}$ has complex codimension $\geqslant 2$, hence by transversality we can always assume that the manifold $S$ does not intersect $E$.
 Indeed, the singular locus of $\pi{\left|{}_{{\phantom{|}\!\!}_{M_{1}}}\right.}$ has codimension 1 in the set $\pi^{-1}(C\cap\Delta)$, where $\Delta$ is the discriminant divisor of $\pi$.
-Since the group structure on the Néron model is continuous, the operation defined in Step 2 can be extended to $S$ considered as a multisection of the Néron model projection to $W_{C}$.
+Since the torsor structure on $M_{2}$ obtained in 3.1 is continuous, the operation defined in Step 2 can be extended to $S$ considered as a multisection of the projection to $W_{C}$.
 This defines a smooth local section of $\pi:\;M_{1}\to C$ over $W_{C}$.
-Step 4: It remains to show that the sections that we constructed locally over $W_{C}\subset C$ agree and define a section of $\pi:\;M_{1}\to C$.
-The possible ambiguity arises from the local choice of the zero section, used in the construction of the Néron model.
-Clearly, it suffices to check that they agree on the smooth locus of $\pi$.
-However, outside of the singularities, the fibers of $\pi$ are tori, equipped with an action of the corresponding abelian group, and the operation (3.1) is independent of the choice of the zero.
-Remark 3.3: This result was conjectured in [KV, Remark 4.2].
+Remark 4.3: This result was conjectured in [KV1, Remark 4.2].
+4.2 Homology classes of sections of abelian fibrations over a line extended to the ambient manifold
+In 4.1 (i) we proved that there exists a smooth section of a Lagrangian fibration restricted to a line ${\mathbb{C}}P^{1}\subset{\mathbb{C}}P^{n}$.
+This section is obtained by taking a multisection of $M_{1}{\>\longrightarrow\>}{\mathbb{C}}P^{1}$ and applying the averaging procedure as explaned in Subsection 4.1, Step 2.
+After an earlier version of this paper was finished, J.
+Kollár contacted us remarking that this procedure might change the homology class of the section (4.2).
+This concern is well founded!
+In the paper [KS], J.
+Kollár and G.
+Saccá studied universal compactified Jacobian $\pi:\;M\to{\mathbb{C}}P^{n}$ of a family of curves on a projective surface.
+This universal compactified Jacobian is a projective variety fibered over a line system of divisors in $S$ (which are curves) with a general fiber over $C\subset S$ identified with the Jacobian $\operatorname{Jac}(C)$.
+A special case of such an abelian fibration is the classical Beauville-Mukai system on a K3 surface (Subsection 1.5).
+Fix a general line $L={\mathbb{C}}P^{1}\subset{\mathbb{C}}P^{n}$, and let $M_{1}:=\pi^{-1}(L)$.
+J.
+Kollár and G.
+Saccá study the restriction map $r:\;H^{2g}(M,{\mathbb{Z}})\to H^{2g}(M_{1},{\mathbb{Z}})$, and show that is not surjective.
+Indeed, for any holomorphic section $s$ of $M_{1}\to L$, such that its fundamental class belongs $\operatorname{im}(r)$, $s$ is necessarily a zero section of $M_{1}\to L$.
+In an email letter, J. Kollár gave the following remarkable and unexpected example.
+Example 4.4: Let $\pi:\;X\to{\mathbb{C}}P^{1}$ be an elliptic fibration on a K3 surface.
+Assume that a special fiber of $\pi$ is a collection of at least 3 rational (-2)-curves, and let $\eta]\in H_{2}(X,{\mathbb{Z}})$ be a homology class of a section.
+Denote by $Z_{1}],...,[Z_{k}]\in H_{2}(X,{\mathbb{Z}})$ the homology classes of the distinct exceptional (-2)-components $Z_{i}$ of a special fiber $F$, and let $n_{1},...,n_{k}$ be a collection of integers.
+Then $A:=\eta+\sum_{i}n_{i}[Z_{i}]$ is a homology class whose intersection with a general fiber is 1.
+Using Thom’s theorem as in Subsection 4.1, Step 2, we can represent $A$ by a smooth 2-dimensional submanifold of $X$.
+The averaging procedure, applied to this situation, produces a section $S$ of the corresponding Néron model, which intersects only one of the irreducible components of $F$, and $A\cap[Z_{i}]=n_{i}[Z_{i}]^{2}+\eta\cap[Z_{i}]$, while $S$, being a section, might intersect only one of $Z_{i}$, with multiplicity 1.
+Remark 4.5: It is not clear how to express the fundamental class of $S$ even if we know $A$ and the numbers $A\cap[Z_{i}]$; this seems to be an interesting question.
+Proof of 4.1 (ii), Step 1: Let $\alpha\in H_{2n}(M,{\mathbb{Z}})$ be an integer homology class which intersects a general fiber of $\pi:M\to CP^{n}$ in 1.
+We represent this homology class by a simplicial chain $A$ with smooth simplices transversal to $\pi$.
+Consider the set $\operatorname{Discr}_{1}\subset{\mathbb{C}}P^{n}$ defined in 3.2, and let $X_{\varepsilon}\subset{\mathbb{C}}P^{n}$ be the complement of its closed $\varepsilon$-neighbourhood.
+We fix a choice of $C$ and $\varepsilon>0$ such that $C\subset X_{\varepsilon}$.
+Recall that Borel-Moore homology is the homology theory dual to cohomology with compact support.
+The homology classes in Borel-Moore homology are represented by locally finite simplicial chains.
+Clearly, $A\cap\pi^{-1}(X_{\varepsilon})$ is homologous to a simplicial chain in Borel-Moore homology, and for an appropriate simplicial decomposition of $X_{\varepsilon}$, we can assume that all simplices of $A_{\varepsilon}$ are mapped to the corresponding simplices of $X_{\varepsilon}$.
+Let ${\mathfrak{X}}_{\varepsilon}$ be a simplicial decomposition of $X_{\varepsilon}$ such that each simplex of $A_{\varepsilon}$ is projected to a certain simplex of ${\mathfrak{X}}_{\varepsilon}$.
+Step 2: By 3.1, the smooth locus of $\pi^{-1}(X_{\varepsilon})$ is an abelian torsor over $X_{\varepsilon}$.
+Since all simplices of $A_{\varepsilon}$ are transversal to the fibers of $\pi$, and $A_{\varepsilon}\cap\pi^{-1}(pt)=1$ in homology, [^1] Here $A_{\varepsilon}\cap\pi^{-1}(pt)=1$ refers to the cap-product in homology.
+the intersection of $A_{\varepsilon}$ and $\pi^{-1}(x)$ is 1, for a general $x$, if counted with multiplicities.
+If we apply the averaging procedure (3.1) using the torsor structure, we obtain that any open simplex $\Delta$ of ${\mathfrak{X}}_{\varepsilon}$ corresponds to an open simplex $s(\Delta)$ obtained by averaging of all simplices of $A$ mapped to $\Delta$.
+It remains to show that this gives a simplicial (hence, continuous) section of $\pi:\;\pi^{-1}(X_{\varepsilon})\to X_{\varepsilon}$.
+Step 3: Let $M_{\varepsilon}:=\pi^{-1}(X_{\varepsilon})$, and $N_{\varepsilon}\subset M_{\varepsilon}$ be the smooth locus. The averaging map
+$\underbrace{N_{\varepsilon}\times_{X_{\varepsilon}}...\times_{X_{\varepsilon}}N_{\varepsilon}}_{\text{$n+1$ times}}\times_{X_{\varepsilon}}\underbrace{N_{\varepsilon}\times_{X_{\varepsilon}}...\times_{X_{\varepsilon}}N_{\varepsilon}}_{\text{$n$ times}}\to N_{\varepsilon}$
+is holomorphic, hence Lipschitz in any compact neighboorhood of $A$.
+All simplices $\Delta$ of the chain $A_{\varepsilon}$ can be chosen smooth, hence Lipschitz.
+Therefore, we can continuously extend each open simplex $s(\Delta)$ to the boundary, obtaining a simplicial chain which gives a simplicial section $\Sigma$ of $\pi$.
+Intersecting this simplicial section with $\pi^{-1}(C)$ where $C$ is a line, we obtain a homology class satisfying the assumptions of 4.1 (ii), as explained in Step 4 below.
+Step 4: The section $\Sigma$ is obtained by taking the homology class $A$ in $H_{2n}(M,{\mathbb{Z}})$, restricting it to the Borel-Moore homology class in $\pi^{-1}(X_{\varepsilon})$, taking the average with respect to the torsor bundle structure (which does not change the Borel-Moore homology class), and restricting the result to the preimage of the curve $C\subset{\mathbb{C}}P^{n}$ by taking the cap-product with $\pi^{-1}(C)$.
+By construction, this is the same as taking a homology class $A\in H_{2n}(M,{\mathbb{Z}})$, and restricting it to $\pi^{-1}(C)$ by taking the cap-product.
 4.3 Constructing a smooth section over a curve (without Néron models)
-We give another proof of 3.1 using the idea communicated to us (after the first draft of this paper was published) by János Kollár.
-This way, we can get rid of the Néron models in 3.1; this also makes it unnecessary to assume that $\pi:\;\pi^{-1}(C)\to C$ has local sections.
+We give another proof of 4.1 (i) using the idea communicated to us (after the first draft of this paper was published) by János Kollár.
+This way, we can get rid of the Néron models in 4.1 (i); this also makes it unnecessary to assume that $\pi:\;\pi^{-1}(C)\to C$ has local sections.
 Theorem 4.6: Let $\pi:\;M\to X$ be a hyperkähler manifold equipped with an Lagrangian fibration, with $X$ smooth (and hence, by Hwang’s theorem, isomorphic to ${\mathbb{C}}P^{n}$), and $C\subset X$ a general line in $X={\mathbb{C}}P^{n}$.
-Assume that the general fiber of the Lagrangian projection $\pi$ is primitive (that is, not divisible) in integer homology.
+Assume that the general fiber of the Lagrangian projection $\pi$ is primitive (that is, not divisible) in the integral homology.
 Then there exists a smooth section $\sigma:\;C{\>\longrightarrow\>}M$.
@@
 Step 1: Let $M_{1}:=\pi^{-1}(C)$.
-Like in the proof of 3.1, Step 2, we can assume that $M_{1}$ contains a smooth submanifold $Z$ such that $[Z]\cap F=1$, where $F=\pi^{-1}(z)$ for a general point $z\in C$.
+Like in the proof of 4.1, Step 2, we can assume that $M_{1}$ contains a smooth submanifold $Z$ such that $Z]\cap F=1$, where $F=\pi^{-1}(z)$ for a general point $z\in C$.
 Let $F_{1}$ be a singular fiber of $\pi$.
 In the first and second step, we show that $Z$ can be chosen in such a way that for some neighbourhood $U\supset F_{1}$, $Z\cap U$ is a disconnected union of holomorphic subvarieties (possibly with an opposite orientation).
-If $F_{1}$ is reduced, this observation is clear: we deform $Z$ in such a way that it intersects $F_{1}$ transversally in its smooth points, and notice that every smooth section of the projection $(z_{1},z_{2},...,z_{n})\mapsto z_{1}$ is homotopic to a holomorphic section near a point $(0,z_{2},...,z_{n})$.
+If $F_{1}$ is reduced, this observation is clear: we deform $Z$ in such a way that it intersects $F_{1}$ transversally in its smooth points, and notice that every smooth section of the projection $(z_{1},z_{2},...,z_{n})\mapsto z_{1}$ is homotopic to a holomorphic section near the point $(0,z_{2},...,z_{n})$.
 Step 2: If $F_{1}$ is not reduced, locally around a smooth point $m$ of its reduction the projection $\pi:\;M_{1}\to C$ can be written as $(z_{1},z_{2},...,z_{n})\mapsto z_{1}^{k}$, where $k$ is the multiplicity of this component of $F_{1}$.
@@
 Step 3: Let $Z_{0}$ be the intersection of $Z$ with the set of all smooth fibers of $\pi:\;M_{1}\to C$.
-These fibers are complex tori, hence we can apply to $Z_{0}$ the averaging procedure defined in 3.1, Step 2, obtaining a section of $\pi:\;M_{1}\to C$ over the set $C_{0}$ of all regular values of this map.
+These fibers are complex tori, hence we can apply to $Z_{0}$ the averaging procedure defined in 3.1, obtaining a section of $\pi:\;M_{1}\to C$ over the set $C_{0}$ of all regular values of this map.
 Since near the critical values $Z$ was a holomorphic multivalued function, and the averaging procedure defines a meromorphic map $M_{1}\times_{\mathbb{C}}M_{1}\times_{C}...\times_{C}M_{1}\to M_{1}$, we obtain a section $S$ of $\pi:\;\pi^{-1}(C)\to C$ which is meromorphic near all singular fibers.
-Since the fibers of $\pi$ are compact, $S$ is holomorphic at each critical value of $\pi:\;\pi^{-1}(C)\to C$ by valuative criterion of properness.
-Therefore this section is smooth on $C$.
+Since the fibers of $\pi$ are compact, $S$ is holomorphic at each critical value of $\pi:\;\pi^{-1}(C)\to C$ by the valuative criterion of properness.
+Therefore, this section is smooth on $C$.
 5 Holography principle for ample rational curves
@@
 [BM], the Barlet space is equipped with a complex structure, which is compatible with the Hausdorff topology.
-Definition 4.1: The Douady space is the moduli of equidimensional dimensional complex subspaces of $M$ (“subspaces” here means that we allow nilpotents in the structure sheaf).
-The Hilbert-Chow morphism is the forgetful map $F:\;{\cal D}{\>\longrightarrow\>}{\cal B}$ from the Douady space to the Barlet space, with multiplicity of each component of $\varphi(\widetilde{Z})$ given by the rank of the structure sheaf of $Z$ in the general point of this component.
+[…]
 Example 5.2: For $\dim Z=0$, the Douady space is the Hilbert space $M^{[n]}$ of 0-dimensional subschemes, and the Hilbert-Chow morphism is the natural projection $M^{[n]}{\>\longrightarrow\>}\operatorname{Sym}^{n}M$.
@@
 This is equivalent to the following definition.
-Definition 4.10:  Let $M$ be an irreducible compact complex variety $M$ bimeromorphic to a compact Kähler manifold.
-Then $M$ is called rationally (chain) connected if for any proper subvariety $E\subset M$, any two points in $M$ can be connected by a sequence of intersecting rational curves which intersect in points of $M\backslash E$.
+[…]
 However, for our purposes, an easier criterion is sufficient.
@@
 Then any holomorphic map $\varphi:\;U_{S}{\>\longrightarrow\>}X$ from $U_{S}$ to a Kähler manifold $X$ can be extended to a meromorphic map $M\dashrightarrow X$.
-Remark 4.13: For 4.3, Kählerness of $X$ is sufficient, but not necessary.
-As seen from the proof, for 4.3 to hold it suffices to establish that all components of the Barlet spaces of curves on $X$ and $M$ are compact.
-This holds true if $X$ admits a Hermitian form $\omega$ which is pluriclosed (i.e., it satisfies $dd^{c}\omega=0$), or such that $-dd^{c}\omega$ is a positive form ([V3]).
-Remark 4.14: It is not difficult to find a counterexample to 4.3 when $X$ is not simply connected and non-Kähler.
-For instance, let $X=\operatorname{Tw}(Y)$ be the twistor space for a 2-dimensional complex torus $Y$ (see e.g.
-[V3]).
-This is an example of a complex manifold with Barlet components of rational curves non-compact (indeed, in [V3] it is shown that all of these components are Stein).
-The universal cover of $\operatorname{Tw}(Y)$ is the twistor space for ${\mathbb{C}}^{2}={\mathbb{H}}$, which can be identified with the total space of ${\cal O}(1)^{2}$ ([JV]).
-Since $\operatorname{Tw}(S^{4})$ is ${\mathbb{C}}P^{3}$ and ${\mathbb{C}}^{2}$ is conformally equivalent to $S^{4}$ without a point, the space $\operatorname{Tw}({\mathbb{C}}^{2})$ is biholomorphic to ${\mathbb{C}}P^{3}\backslash{\mathbb{C}}P^{1}$ (see [JV] for another argument leading to this conclusion).
-Then an open neighbourhood of ${\mathbb{C}}P^{1}\subset{\mathbb{C}}P^{3}$ can be embedded into $\operatorname{Tw}(Y)$.
-However, if this embedding can be extended to the whole ${\mathbb{C}}P^{3}$, the space $\operatorname{Tw}(Y)$ would be Moishezon, being an image of a proper holomorphic map from a Moishezon space (see e.g.
-[F]).
-This is impossible, because the algebraic dimension of any twistor space of a compact hyperkähler manifold is one ([V2]).
+Remark 5.13: For 5.3, Kählerness of $X$ is sufficient, but not necessary.
+As seen from the proof, for 5.3 to hold it suffices to establish that all components of the Barlet spaces of curves on $X$ and $M$ are compact.
+This holds true if $X$ admits a Hermitian form $\omega$ which is pluriclosed (i.e., it satisfies $dd^{c}\omega=0$), or such that $dd^{c}\omega$ is a positive form ([V3]).
+[…]
 We expect that a similar counterexample would exist when $X$ is non-Kähler but simply connected.
-When $X$ is Moishezon, 4.3 is implied by the following theorem (see also [V2, Theorem 3.4]; if we deal with formal schemes instead of complex manifolds, this result was established in [Har] and in [HM, Hir]).
+When $X$ is Moishezon, 5.3 is implied by the following theorem (see also [V2, Theorem 3.4]; if we deal with formal schemes instead of complex manifolds, this result was established in [Har] and in [HM, Hir]).
 The following result requires the manifold $M$ to be homogeneous and rational; for non-homogeneous case, see [V2, Theorem 3.4].
@@
 Note that “algebraic map” in [Chow, Theorem 2] refers to multi-valued maps given by an algebraic correspondence, which is shown to be rational in [Chow, Theorem 3].
-Proof of 4.3: The main idea of the proof of 4.3 is the following observation.
+Proof of 5.3: The main idea of the proof of 5.3 is the following observation.
 Consider an irreducible complex subvariety $B_{0}$ in the Barlet space of curves on $X$ containing $n\varphi(S^{\prime})$, where $n$ is a positive integer (representing multiplicity) and $S^{\prime}\subset U_{S}$ a general deformation of $S$ in $U_{S}$.
@@
 Assume that $L_{0}$ is rationally connected.
-Then $L_{0}$ is Moishezon ([C1]) and its normalization is simply connected ([Deb, Corollary 2.32]; see also [C2, KMM]), hence 4.3 can be applied, showing that the map $\varphi:\;U_{S}{\>\longrightarrow\>}L_{0}$ can be extended to a meromorphic map $M\dashrightarrow L_{0}\subset X$.
+Then $L_{0}$ is Moishezon ([C1]) and its normalization is simply connected ([Deb, Corollary 2.32]; see also [C2, KMM]), hence 5.3 can be applied, showing that the map $\varphi:\;U_{S}{\>\longrightarrow\>}L_{0}$ can be extended to a meromorphic map $M\dashrightarrow L_{0}\subset X$.
 Step 1: We take $n=3$.
@@
 To finish the proof, it remains to show that $L_{0}$ is rationally connected.
-We will prove that two generic points of $L_{0}$ are connected by a smooth curve which belongs to $B_{0}$; by 4.2, this implies the rational connectedness of $L_{0}$.
+We will prove that two generic points of $L_{0}$ are connected by a smooth curve which belongs to $B_{0}$; by 5.2, this implies the rational connectedness of $L_{0}$.
 Step 2: Since $S$ is ample, its normal bundle $NS$ is a direct sum of ${\cal O}(i)$ with $i\geqslant 1$.
@@
 Moreover, we can perform this deformational smoothing in such a way that $S_{3}$ contains $a$ and $b$.
-This is the idea used to prove 4.2.
+This is the idea used to prove 5.2.
 Start from $a,b\in U_{S}$ generic and sufficiently close to $S$, and let $S_{1}$ connect $a$ to $x$, $S_{2}$ connect $b$ to $x$, and $S_{3}$ connect $z\in S$ to $t\in S_{1}$.
@@
 This was done in Step 2.
-As shown in Step 1, this implies that $L_{0}$ is rationally connected, and finishes the proof of 4.3.
+As shown in Step 1, this implies that $L_{0}$ is rationally connected, and finishes the proof of 5.3.
 6 Meromorphic sections via Dolbeault currents
@@
 Definition 6.1: Let $\pi:\;M{\>\longrightarrow\>}X$ be a Lagrangian fibration, and $\sigma:\;X{\>\longrightarrow\>}M$ a section.
-By [BDV, Proposition 2.1] (see also 5.2 below), $\sigma^{*}\Omega$ is a $(2,0)+(1,1)$ closed form on $X$.
+By [BDV, Proposition 2.1] (see also 6.2 below), $\sigma^{*}\Omega$ is a $(2,0)+(1,1)$ closed form on $X$.
 Assume that $H^{2}(X)$ is generated by a closed $(1,1)$-form $\omega$.
@@
 Let $\alpha$ be a 1-form which satisfies $d\alpha=\sigma^{*}\Omega-\lambda\omega$.
-Then ¯$\overline{\partial}\alpha^{0,1}=0$.
+Then $\overline{\partial}\alpha^{0,1}=0$.
 The Dolbeaut class of $\sigma$ is the class of $\alpha^{0,1}$ in $H^{1}(\Lambda^{0,*}X,\overline{\partial})=H^{1}(X,{\cal O}_{X})$.
 Proposition 6.2: Let $\pi:\;M{\>\longrightarrow\>}{\mathbb{C}}P^{n}$ be a Lagrangian fibration, $C\subset{\mathbb{C}}P^{n}$ a smooth curve, $C_{U}$ its neighbourhood and $\sigma_{U}:\;C_{U}{\>\longrightarrow\>}M$ a section.
-Assume that the Dolbeault class $[\alpha^{0,1}]\in H^{1}(C_{U},{\cal O}_{C_{U}})$ of $\sigma_{U}$ vanishes.
+Assume that the Dolbeault class $\alpha^{0,1}]\in H^{1}(C_{U},{\cal O}_{C_{U}})$ of $\sigma_{U}$ vanishes.
 Then $\sigma^{*}_{U}\Omega$ restricted to a smaller neighbourhood of $C$ can be extended to a closed $(2,0)+(1,1)$-form on ${\mathbb{C}}P^{n}$.
@@
 Let $\rho$ be closed $(2,0)+(1,1)$-form on $U$, and $\rho=d\alpha+\lambda\omega$.
-Then the $(0,1)$-Dolbeault class of $\rho$ is $[\alpha^{0,1}]\in H^{0,1}(M)$.
+Then the $(0,1)$-Dolbeault class of $\rho$ is $\alpha^{0,1}]\in H^{0,1}(M)$.
 Remark 6.4: By definition, the Dolbeault class of $\sigma_{U}$ is equal to the (0,1)-Dolbeault class of $\sigma_{U}^{*}\Omega$.
 In the conclusion of this section (Subsection 6.3), we prove the following theorem:
-Theorem 5.5: Let $\pi:\;M{\>\longrightarrow\>}{\mathbb{C}}P^{n}$ be a Lagrangian fibration, $C\subset{\mathbb{C}}P^{n}$ a general line, let $C_{U}$ be a neighbourhood of $C$, and $\sigma_{U}:\;C_{U}{\>\longrightarrow\>}M$ a section.
+Theorem 6.5: Let $\pi:\;M{\>\longrightarrow\>}{\mathbb{C}}P^{n}$ be a Lagrangian fibration, $C\subset{\mathbb{C}}P^{n}$ a general line, let $C_{U}$ be a neighbourhood of $C$, and $\sigma_{U}:\;C_{U}{\>\longrightarrow\>}M$ a smooth section which satisfies 4.1 (ii).
 Then the Dolbeault class of $\sigma_{U}$ vanishes.
-Proof: 5.3.
+Proof: 6.3.
 This theorem completes the last step of the proof of 1.1, the main result of this paper. Before we prove this theorem, we need to make a digression to the land of currents.
@@
 Definition 6.11: On a complex manifold, $(p,q)$-currents are $(p,q)$-forms with coefficients in generalized functions.
-Remark 5.12: In the literature, this is sometimes called “$(n-p,n-q)$-currents”; we follow the conventions from [Dem].
+Remark 6.12: In the literature, this is sometimes called “(np,nq)(n-p,n-q)-currents”; we follow the conventions from [Dem].
 Claim 6.13: De Rham differential is continuous on currents, and the Poincaré lemma holds for currents, as well as the Poincaré-Dolbeault-Grothendieck lemma.
@@
 Define the pushforward $f_{*}\alpha$ using $\langle f_{*}\alpha,\tau\rangle:=\langle\alpha,f^{*}\tau\rangle$, where $\tau$ is any test-form.
-Then $f_{*}\alpha$ is a $(p-k,q-k)$-current.
+Then $f_{*}\alpha$ is a (pk,qk)(p-k,q-k)-current.
 One should think of $f_{*}$ as of the fiberwise integration.
@@
 Let $\dim_{\mathbb{C}}M=2n$.
-To prove that $\pi_{*}(\Omega\wedge[Z])$ is of type $(2,0)+(1,1)$ it suffices to show that $\pi_{*}(\Omega\wedge[Z])\wedge\tau=0$ for any $(n,n-2)$-form $\tau$.
+To prove that $\pi_{*}(\Omega\wedge[Z])$ is of type $(2,0)+(1,1)$ it suffices to show that $\pi_{*}(\Omega\wedge[Z])\wedge\tau=0$ for any (n,n2)(n,n-2)-form $\tau$.
 This would follow if we prove that $\pi^{*}\tau\wedge\Omega=0$.
@@
 6.3 Vanishing of the cohomology class of the Dolbeault current
-Theorem 5.18:  Let $\pi:\;M{\>\longrightarrow\>}{\mathbb{C}}P^{n}$ be a Lagrangian fibration, $C\subset{\mathbb{C}}P^{n}$ a general line, $C_{U}$ a neighbourhood of $C$, and $\sigma_{U}:\;C_{U}{\>\longrightarrow\>}M$ a section constructed in 3.1.
+Theorem 6.18: Let $\pi:\;M{\>\longrightarrow\>}{\mathbb{C}}P^{n}$ be a Lagrangian fibration, $C\subset{\mathbb{C}}P^{n}$ a general line, $C_{U}$ a neighbourhood of $C$, and $\sigma_{U}:\;C_{U}{\>\longrightarrow\>}M$ a smooth section which satisfies 4.1 (ii).
 Then the Dolbeault class of $\sigma_{U}$ vanishes.
@@
 This is done as follows: we split $Z$ onto smooth pieces, restrict $\alpha$ to each, transport to $X$ with the appropriate sign and sum together, and then take the $\overline{\partial}$-class of the (0,1)-part of resulting current.
-This is equivalent to taking the current $\pi_{*}(\alpha\wedge[Z])$, using 5.2 to show that its (0,1)-part is $\overline{\partial}$-closed, and taking its $\overline{\partial}$-class.
+This is equivalent to taking the current $\pi_{*}(\alpha\wedge[Z])$, using 6.2 to show that its (0,1)-part is $\overline{\partial}$-closed, and taking its $\overline{\partial}$-class.
 Step 2: With this definition, we can extend the notion of a Dolbeault class to any multisection of a Lagrangian fibration $\pi:\;M{\>\longrightarrow\>}{\mathbb{C}}P^{n}$, and, indeed, to any piecewise smooth homology chain $u\in H^{2n}(M,{\mathbb{Z}})$, where $\dim_{\mathbb{C}}M=2n$.
@@
 Then the Dolbeault class of $\sigma$ is equal to the $(0,1)$-Dolbeault class of the current $d(\pi_{*}(\Omega\wedge s))$, but the (0,1)-Dolbeault class of an exact current vanishes.
-Step 3: As shown in 3.1 (step 1 of the proof), the section $\sigma_{C}:\;C{\>\longrightarrow\>}M$ is obtained as a representative of an integral homology class on $M$ intersected with $\pi^{-1}(C)$, hence its homology class can be extended to $M$.
+Step 3: Since the section $\sigma_{C}$ satisfies 4.1 (ii), its homology class is obtained as a cap product of $\pi^{-1}(C)$ with an integer homology class $\alpha\in H_{2n}(M,{\mathbb{Z}})$.
 This implies that the Dolbeault class of $\sigma_{U}$ vanishes; indeed, a restriction of any class in $H^{0,1}({\mathbb{C}}P^{n})$ to $C_{U}$ vanishes, because $H^{0,1}({\mathbb{C}}P^{n})=0$.
-Combining this theorem with 4.3, we obtain the main result of this paper.
+Combining this theorem with 5.3, we obtain the main result of this paper.
 Theorem 6.19: Let $\pi:\;M{\>\longrightarrow\>}{\mathbb{C}}P^{n}$ be a Lagrangian fibration on a hyperkähler manifold.
@@
 Step 1: In [SV2], it was shown that any degenerate twistor deformation of a hyperkähler manifold is Kähler.
-Let $C\subset{\mathbb{C}}P^{n}$ be a general line in the base, and $\sigma:\;C{\>\longrightarrow\>}M$ a smooth section constructed in 3.1.
+Let $C\subset{\mathbb{C}}P^{n}$ be a general line in the base, and $\sigma:\;C{\>\longrightarrow\>}M$ a smooth section constructed in 4.1.
 We extend it to an open neighbourhood $C_{U}$ of $C$, obtaining a map $\sigma_{U}:\;C_{U}{\>\longrightarrow\>}M_{U}$, where $M_{U}:=\pi^{-1}(C_{U})$.
 We apply 1.2 to construct a degenerate twistor deformation $M^{\prime}_{U}$ of $M_{U}$ such that the section $\sigma_{U}:\;C_{U}{\>\longrightarrow\>}M_{U}^{\prime}$ is holomorphic.
-If this degenerate twistor deformation can be extended from $M_{U}^{\prime}{\>\longrightarrow\>}C_{U}$ to $M^{\prime}{\>\longrightarrow\>}{\mathbb{C}}P^{n}$, then 4.3, applied to the holomorphic map $C_{U}{\>\longrightarrow\>}M^{\prime}$, gives a rational section of $\pi$.
+If this degenerate twistor deformation can be extended from $M_{U}^{\prime}{\>\longrightarrow\>}C_{U}$ to $M^{\prime}{\>\longrightarrow\>}{\mathbb{C}}P^{n}$, then 5.3, applied to the holomorphic map $C_{U}{\>\longrightarrow\>}M^{\prime}$, gives a rational section of $\pi$.
 Step 2: Let $\eta_{U}:=\sigma_{U}^{*}(\Omega)$.
-By 5.2, $\eta_{U}$ is of type $(2,0)+(1,1)$.
+By 6.2, $\eta_{U}$ is of type $(2,0)+(1,1)$.
 To obtain $M^{\prime}$, we have to extend $\eta_{U}$ to a closed form of type $(2,0)+(1,1)$ on ${\mathbb{C}}P^{n}$.
@@
 [^1] It is possible to see that this group is infinite-dimensional.
-Step 3: In 5.3 we proved that the obstruction class in $H^{1}(C_{U},{\cal O})$ (i.e., “the Dolbeault class of the section $\sigma_{U}$”) vanishes.
-Now, 5.1 implies that the degenerate twistor deformation $M_{U}^{\prime}$ can be extended to $M^{\prime}{\>\longrightarrow\>}{\mathbb{C}}P^{n}$.
+Step 3: In 6.3 we proved that the obstruction class in $H^{1}(C_{U},{\cal O})$ (i.e., “the Dolbeault class of the section $\sigma_{U}$”) vanishes.
+Now, 6.1 implies that the degenerate twistor deformation $M_{U}^{\prime}$ can be extended to $M^{\prime}{\>\longrightarrow\>}{\mathbb{C}}P^{n}$.
 Acknowledgments.